### v1

# T-MAE: Temporal Masked Autoencoders for Point Cloud Representation Learning

Abstract
The scarcity of annotated data in outdoor point cloud segmentation poses a significant obstacle in harnessing the modeling capabilities of advanced networks like transformers. Consequently, scholars have been actively investigating efficacious self-supervised pre-training strategies, e.g. contrasting learning and reconstruction-based pretext tasks. Nevertheless, temporal information, which is inherent in the LiDAR point cloud sequence, is consistently disregarded. To better utilize this property, we propose an effective pre-training strategy, namely Temporal Masked Auto-Encoders (T-MAE), which takes as input temporally adjacent frames and learns temporal dependency. A SiamWCA backbone, containing a Siamese encoder and a window-based cross-attention (WCA) module, is established for the two-frame input. Taking into account that the motion of an ego-vehicle alters the illumination angles of the same instance, temporal modeling also serves as a robust and natural data augmentation, enhancing the comprehension of target objects. Moreover, instead of utilizing consecutive frames, it is more cost-effective and powerful by using distant historical frames. SiamWCA is a powerful architecture but heavily relies on annotated data. With our T-MAE pre-training strategy, we achieve the best performance on the Waymo dataset among self-supervised learning methods. Comprehensive experiments are conducted to validate all components of our proposal. Upon acceptance, the source code will be made accessible.
1 Introduction
Figure 1: T-MAE performance on Waymo dataset [39]. Our T-MAE pre-training achieves substantially better results than random initialization. It also achieve better results than the SOTA SSL method MV-JAR [48] with significantly fewer iterations. Each point triplet corresponds to a specific fine-tuned model variant and shows the performance difference with two arrows. From left to right, the triplets correspond to a model fine-tuned with 5%, 10% and 20% data.
In the domains of language and vision, researchers have observed that as models become more complex, the available amount of labeled data is often insufficient to adequately train huge models [11, 18] e.g. Vision Transformer (ViT) [12]. Consequently, there is a growing interest in exploring self-supervised learning (SSL) approaches as a potential solution to overcome this limitation. SSL serves as a pre-training technique with unlabeled data, accelerating the convergence of the models and improving their performance for downstream tasks [16, 6, 9]. The same challenge extends to the domain of point clouds, where annotations are more costly and time-consuming to obtain [46, 14]. For instance, a mere 10% and 0.8% of sweeps are annotated in the nuScene [2] and ONCE [29] datasets. This challenge makes pre-training for point cloud understanding a non-trivial endeavor.
Prior works mainly focus on synthetic and isolated objects [50, 56, 54, 35, 27] and indoor scene understanding [19, 24, 49, 36]. Transferring these methods to outdoor LiDAR points is challenging due to their sparsity and dynamic environmental conditions. At present, most self-supervised methods used for understanding point clouds in autonomous driving rely on contrastive learning [47, 53, 32, 34, 33]. These approaches model the similarity and dissimilarity between entities, such as segments [33, 45] and/or points [47]. In the wake of masked image modeling as a pretext task [18], efforts have also been devoted to the reconstruction of masked points [31, 41, 52, 48]. The main idea is randomly masking points or voxels and urging the network to infer the coordinates of points [52] and/or voxels [48] or other properties, e.g. occupancy [31, 1] and curvature [41]. Nevertheless, these methodologies often operate within the confines of a single-frame scenario, disregarding the fact that LiDAR data is typically acquired on a frame-by-frame basis. In other words, the valuable semantic information that is introduced by temporally adjacent frames is often not exploited.
Several methods attempt to leverage temporal information [21, 45, 33, 25] by incorporating multi-frame input during the self-supervised phase but their core concepts remain grounded in contrastive learning. Specifically, the scans captured at different times are treated as augmented samples of the same scene, without including temporal correspondence into the modeling procedure.
Therefore, we propose a new self-supervised method, namely T-MAE, to exploit the accumulated observations. The key idea of T-MAE is to learn a powerful representation for point cloud scans and integrate features from the past for present use. During the pre-training stage, the current scan is voxelized with a high masking ratio, while the previous scan is fed entirely to the encoder. Then, the reconstruction pretext task is performed by incorporating three types of information: voxel embeddings of the past scan, visible voxel embeddings of the current scan, and the position of masked voxels. In this way, the WCA module learns to incorporate historical information into the current frame using unlabeled data. The proposed T-MAE pre-training strategy endows the network with both a powerful representation for sparse point clouds and the capacity to strengthen the present by learning from the past. As shown in Fig. 1, the proposed T-MAE achieves higher overall and pedestrian-specific mAPH than the randomly initialized baseline, namely the same model but trained from scratch. Moreover, T-MAE also outperforms state-of-the-art MV-JAR [48] with over 1.6×1.6$\times$1.6 × fewer training iterations. Our contributions are summarized as follows:
  • We propose T-MAE, a novel and effective SSL approach for sparse point clouds, that learns temporal modeling in the process of reconstructing masked points.
  • We design a SiamWCA backbone, containing a Siamese encoder and a window-based sparse cross-attention (WCA) module, to incorporate historical information.
  • Rather than concatenating multiple consecutive point cloud scans, our approach demonstrates that incorporating more distant historical frames provides greater advantages.
  • Our experiments demonstrate the efficacy of T-MAE by attaining impressive results on the Waymo Open dataset. T-MAE achieves 55.71 mAPH for pedestrians with only 5% of the labeled data and surpasses the SOTA SSL approaches by a substantial margin across all levels of fine-tuning data amount.
2 Related Work
2.1 Static Self-supervised Learning
SSL for point clouds is a burgeoning field due to the scarcity of annotations. In general, the pipeline of SSL consists of two phases. During the initial stage, namely pre-training with unlabeled data, the network is trained using a pretext task. Subsequently, in downstream tasks such as segmentation and detection, the pre-trained weights are loaded to the backbone, and the backbone is attached with task-specific heads. The resulting model is further fine-tuned using annotated data.
The initial research efforts are primarily directed towards the exploration of contrastive learning. The underlying hypothesis is that an image or 3D scene demonstrates feature equivalence even after undergoing different transformations [9, 5]. One of the key challenges in the point cloud domain is establishing correspondences across scenes. PointContrast [47] and DepthContrast [57] track the points while performing different transformations. GCC-3D [25] exploit sequential information to obtain pseudo instances and then perform contrasting on the instance level. SegContrast [32] firstly obtains segments by an unsupervised clustering and then contrasting segments between two transformed views. These methods allow the network to learn equivalence with regard to geometric transformations. However, contrastive learning-based SSL approaches usually suffer from careful tuning of hyperparameters and complicated pre- or post-processing to find correspondences.
The focus has shifted to reconstructing masked points, following the success of masked image modeling [18, 44]. The masking strategies remain consistent, i.e. voxelize a point cloud into cubes followed by random masking of these cubes. The reconstruction targets vary across papers. OccupancyMAE [31] classifies voxels to whether they are occupied. Geo-MAE [41] infers occupancy as well as the normal and the curvature of each masked voxel. The GD-MAE’s pretext task is to reconstruct a fixed number of points for masked voxels [52]. MV-JAR [48] partitions the masked Voxel into two categories, necessitating the network to either predict the coordinates of the voxels or produce the points themselves. In this paper, we employ the same masking strategy as GD-MAE [52] but expand the knowledge source to include both the current visible voxels and voxels from a previous frame as a reference.
2.2 Temporal Self-supervised Learning
There are many SSL methods designed to address temporal or spatiotemporal tasks. In video-related tasks, such as MAE-ST [15], a random patch masking strategy is employed over consecutive frames. The model is tasked with recovering these masked patches while considering information from neighboring frames. VideoMAE [42] retains the same reconstruction pipeline as the pretext task but exploits a tube masking strategy. SiameseMAE [17] learns object-centered representations with the help of cross-attention layers and an asymmetrical masking technique on consecutive frames. There are also content-based masking strategies, e.g. motion-guided masking [20, 30]. The primary concept underlying these methods is acquiring an understanding of temporal dependency across frames through the process of reconstructing patches with reference to consecutive frames.
While this concept is effectively used in video understanding, it is rarely applied to the learning of sparse point cloud representations. This is because point clouds are typically handled on a frame-by-frame basis rather than being regarded as a temporal sequence. For instance, when using temporally adjacent scans as input during self-supervised learning, both STSSL [45] and TARL [33] use HDBSCAN [3] to obtain segments and apply self-supervision by minimizing the feature distance between the same segments or points between neighbor frames. While they achieve impressive results, two issues persist 1) The use of elaborate and time-consuming pre-processing methods to obtain segments. 2) A primary focus on object consistency across frames rather than understanding object motion.
Unlike other approaches, this paper aims to reduce the reliance on complex preprocessing techniques and emphasizes the significance of enabling the model to establish temporal correspondence through self-supervised learning from unlabeled data.
3 Method
Figure 2: Overview of the proposed T-MAE pre-training. Two frames are sampled from a sequence of point clouds according to temporal batch-based strategy. The prior scan $\mathcal{P}^{t+t_{1}}$ is voxelized and directly fed into the Siamese windowed cross-attention (SiamWCA) backbone. The current scan $\mathcal{P}^{t+t_{2}}$ undergoes a masking process before fed into the backbone. Furthermore, pillar-wise tokens are computed by a Siamese hierarchical encoder. The two-way gray arrow indicates weight sharing. The WCA module takes as input the full tokens of the previous frame and the partial observation of the current frame and creates enhanced tokens. In the diffusion process, masked tokens are updated by their neighbor tokens from WCA by a group of convolution layers. Subsequently, a reconstruction head recovers masked points. During fine-tuning, the full SiamWCA backbone is initialized with pre-trained weights obtained from the pre-training phase.
In this section, previous works are briefly discussed in Sec. 3.1. Then, the proposed T-MAE pre-training strategy is demonstrated in Sec. 3.2. Eventually, the window-based sparse cross-attention (WCA) module is elaborated in Sec. 3.3.
3.1 Preliminaries
Pillar-based representation and sparse regional self-attention.
The pillar-based representation is an efficient representation introduced in PointPillars [23]. It divides 3D points into infinite-height voxels and computes pillar-wise features. These features can be treated as a pseudo-image in a bird’s eye view. On top of the pillar-based representation, SST [13] proposes a sparse regional self-attention (SRA) module to address a challenge in applying ViT to sparse LiDAR points. SST divides the pillars into non-overlapping windows with a fixed size and then applies self-attention to the non-empty pillars within the same window. The receptive field of each pillar is expanded through a window shift operation. T-MAE and many recent works [52, 48, 40] are built up on the basic block SRA.
Temporal batch.
A sweep of a point cloud at time $t$ is denoted by $\mathcal{P}^{t}=\{(p^{t}_{k})\}_{k=1}^{K}$, where $p_{k}$ is a 3D point $p_{k}\in\mathbb{R}^{3}$. Moreover, a sequence of point clouds is an ordered set of $\mathcal{P}^{t}$, which is denoted by $\mathcal{P}=\{\mathcal{P}^{1},\mathcal{P}^{2},\ldots,\mathcal{P}^{T}\}$. During pre-training and fine-tuning, two scans are needed feed the Siamese encoder. Due to the high sampling frequency of the LiDAR system, two consecutive frames usually contain redundant duplicate information, which is verified in Sec. 5. Conversely, when there is a substantial time gap between two frames, their overlap may be limited, rendering the historical information less useful. Therefore, we follow TARL [33] to introduce the concept of a temporal batch so that the interval of two frames is constrained to an appropriate range. Specifically, we sample a batch of consecutive frames $\mathcal{B}^{t}=\{\mathcal{P}^{t+1},\mathcal{P}^{t+2},\ldots,\mathcal{P}^{t+n}\}$ from a sequence $\mathcal{P}$. Then, two frames, $\mathcal{P}^{t+t_{1}}$ and $\mathcal{P}^{t+t_{2}}$ are sampled from the first and last one-third of the batch. More specifically, $0\leq t_{1}<\frac{n}{3}$ and $\frac{2n}{3}\leq t_{2}$.
3.2 Temporal Masked Autoencoder Pre-training
Inspired by the SiamMAE [17] used in video understanding, we develop Temporal Masked Auto-Encoders, namely T-MAE, for self-supervised learning on LiDAR points. As shown in Fig. 2, the core idea is to reconstruct the present frame based on the historical and partial observations of the current frame. In this way, the network is compelled to learn a powerful sparse representation as well as the capacity to effectively model motion. After the pre-training, the weights of the SiamWCA backbone are retained for the downstream tasks. In the subsequent subsections, we provide a detailed elaboration of these steps.
Single-frame baseline.
To perform the reconstruction pretext task, we follow the concept of the state-of-the-art SSL method GD-MAE [52] which operates on a single frame. The pipeline encompasses several key stages, including voxelization, masking, encoding, decoding, and reconstruction. More specifically, the initial step voxelizes a given point cloud by two linear layers and a voxel-wise average pooling. The voxel representation can be considered a pillar-based representation, owing to its infinite height. Next, masking is applied in a pillar-wise manner. Subsequently, the remaining visible pillars serve as input for the sparse pyramid transformer to generate high-level tokens. Next, the sparse tokens are transformed into a dense tensor by means of zero-padding. Subsequently, a series of dense convolutional layers are utilized to dilate the tokens of unmasked voxels to vacant regions.
$t$ ${t-3}\ \sim\ t$
Refer to caption Refer to caption
(a) A stationary vehicle.
Refer to caption Refer to caption (b) A moving vehicle.
Figure 3: Comparison between single and four frame concatenation. While simple frame concatenation generally improves point density and detection rates, it can introduce spurious points in non-static scene parts that may degrade the detection performance. Since we combine consecutive frames via learned cross-attention, our approach is less affected by this problem. The blue bounding boxes indicate the ground truth for the current frame.
Figure 4: Window-based sparse cross-attention (WCA) module. Given the input tokens from both the previous and current frames, a joint region grouping is conducted to obtain a window partition. A sparse regional cross-attention (SRCA) is performed individually in each window to integrate the historical information to the middle tokens for the current frame. In other words, the tokens from two frames but with the same colors are attending to each other. For simplicity, the information flow is only illustrated for the green tokens. The grouping and cross-attention are performed once more with the shifted window partition. The red dot (Refer to caption) indicates the ego-vehicle driving towards the right. The box with diagonal stripes (Refer to caption) represents an object moving towards the left. The observed scene remains mostly static except for a pillar at bottom-left transitions from occupied to vacant due to occlusion and another pillar which does the opposite. Best viewed in color and high-resolution.
SiamWCA.
Up to this point, a common approach has been to employ a single-frame reconstruction pretext task for LiDAR points. However, incorporating historical frames poses a non-trivial challenge. One straightforward approach is to concatenate two point clouds after aligning them with ego-poses, similar to the three-frame variant of SST [13]. However, as shown in Fig. 3, while concatenation helps to identify static objects effectively, it introduces challenges for detectors when dealing with moving objects. We provide quantitative results in Tab. 2. To address this problem, we suggest learning from past data in a feature space.
Given two frames, the previous frame is transformed to the coordinate system of the later frame by their ego-poses. Pose information is usually readily available in relevant datasets [39, 2, 29] or easily obtained by means of GPS/IMU, odometry approaches [8], structure from motion (SfM) algorithms [37, 38], or SLAM systems [10]. As shown in Fig. 2, the full observation of the previous frame is given to the network by passing it through voxelization and encoder. On the contrary, the later frame goes through voxelization, masking, and encoder. The encoders for the two frames share weights to ensure that the features are constrained in the same space.
By utilizing the Siamese encoder, tokens are obtained for both frames. These tokens serve as input to the WCA module, which facilitates the interaction between historical and current tokens and is outlined in Sec. 3.3. Once the current tokens are augmented with historical information, these sparse tokens are placed back into the x-y plane to form a dense tensor where vacant pillars are padded with zeros. Subsequently, a group of $3\times 3$ convolutional layers is utilized to process the tensor, which diffuses the features of visible pillars. Eventually, the obtained tokens of masked pillars are employed for point reconstruction. The Chamfer distance is computed as the loss function between the reconstructed points and the ground-truth points. Note that the number of points per pillar varies drastically. Thus, a fixed number of points $K^{GT}$ are randomly sampled as the target for reconstruction.
In conclusion, SiamWCA, containing a Siamese encoder and a WCA module, computes the pillar-based representation for both frames and integrates the historical one to the present one for reconstruction. Unlike the single-frame baseline where only the encoder is reused for downstream tasks, the weights of SiamWCA are fully retained for downstream tasks.
3.3 Window-based Sparse Cross-Attention Module
When provided with input tokens from two frames, there are several methods for interaction. One intuitive approach is to apply a standard self-attention layer to the concatenation of all tokens from both frames. This approach significantly increases GPU memory requirements. The vanilla Transformer block [43], which is combined with a cross-attention layer and a self-attention layer, can also be adopted. This combination, however, is demonstrated to be suboptimal as outlined in Sec. 4.3. In due course, a single cross-attention layer is evaluated and exhibits superior performance compared to others.
However, conventional global attention is not affordable in 3D space due to the expensive computational overhead imposed by the unavoidable high resolution. Therefore, the use of window-based attention is vital. The WCA module is illustrated in Fig. 4. The simulation involves an ego-vehicle moving to the right and another object moving in the opposite direction. The scene remains mostly static, but a pillar at bottom-left transitions from occupied to vacant due to occlusion.
Joint Regional Grouping
organizes the 3D spaces into non-overlapping windows. Hence, the following attention mechanism only performs with tokens from the same local region. The tokens from two frames are jointly considered so that the local windows across the two frames are corresponding. As shown in Fig. 4, all tokens within the same physical window share the same color, indicating that they are assigned to the same group for mutual attention.
Sparse Regional Cross-Attention
(SRCA) is essentially a cross-attention layer where the query comes from the current frame and the key-value comes from the previous frame. For a group of tokens $\mathcal{F}^{t_{1}},\mathcal{F}^{t_{2}}$ and their corresponding spatial coordinates $\mathcal{I}^{t_{1}},\mathcal{I}^{t_{2}}$, the query $\mathcal{Q}$, key $\mathcal{K}$, and value $\mathcal{V}$ are calculated as follows
$\mathcal{Q}$ $=\mathcal{F}^{t_{2}}+\mathbf{PE}(\mathcal{I}^{t_{2}})$ (1)
$\mathcal{K}$ $=\mathcal{F}^{t_{1}}+\mathbf{PE}(\mathcal{I}^{t_{1}})$ (2)
$\mathcal{V}$ $=\mathcal{F}^{t_{1}}$ (3)
where $\mathbf{PE}(\cdot)$ represents the absolute positional encoding function used in [4]. SRCA follows conventional cross-attention and is defined as follows
^$\hat{\mathcal{F}}^{t_{2}}$ $=\mathbf{MCA}(\mathcal{Q},\mathcal{K},\mathcal{V})$ (4)
~$\widetilde{\mathcal{F}}^{t_{2}}$ $=\mathbf{LN}\left(\mathbf{MLP}\big{(}\mathbf{LN}(\hat{\mathcal{F}}^{t_{2}})\big{)}+\hat{\mathcal{F}}^{t_{2}}\right)$ (5)
where $\mathbf{MCA}(\cdot,\cdot,\cdot)$ indicates the Multi-head Cross-Attention, and $\mathbf{LN}(\cdot)$ stands for Layer Normalization. While the core concept is similar to cross-attention, this implementation provides a more efficient solution when working with sparse pillars.
Repeated Operations with Region Shift.
The joint regional grouping and SRCA are performed once more after the regions are shifted. We follow the implementation of SST [13] to achieve region shift. Formally, suppose that the window size is $(l_{x},l_{y},l_{z})$, then the original windows are moved by $(l_{x}/2,l_{y}/2,l_{z}/2)$. Meanwhile, the tokens are re-grouped by the joint regional grouping, as illustrated in “Shifted Window Partition” of Fig. 4. Note that, while performing the SRCA #2, the tokens ~$\widetilde{\mathcal{F}}^{t_{2}}$ of the current frame are updated by SRCA #1 while the tokens $\mathcal{F}^{t_{1}}$ remain unchanged.
4 Experiments
4.1 Dataset and Implementation
Waymo Open dataset [39] is the currently the largest public dataset with LiDAR points for autonomous driving. It comprises 798 sequences for training and 202 sequences for validation. Each sequence contains roughly 200 frames with pose information by which points can be transformed into the world coordinate system. For 3D detection, an evaluation protocol is provided for calculating the average precision (AP) and the average precision weighted by heading (APH). Moreover, the evaluation includes two difficulty levels wherein bounding boxes containing over 5 points are regarded as Level 1 while Level 2 indicates all bounding boxes. We adopt mean AP and mean APH at Level 2 as main evaluation metrics.
Implementation Details. We implement our approach based on the codebase of OpenPCDet [^1] https://github.com/open-mmlab/OpenPCDet. We keep the configuration of the sparse pyramid transformer consistent with its official implementation of GD-MAE [52]. The window size of the WCA modules is set to $(8,8,1)$. The detection ranges for the Waymo dataset are restricted between $(-74.88m,-74.88m,-2m)$ and $(74.88m,74.88m,4m)$. The pillar size is $(0.32m,0.32m,6m)$. The size of the temporal batch is set to $n=6$ for both pre-training and finetuning. Three data augmentation techniques, i.e. random flipping, scaling, and rotation, are applied to both frames during pre-training and finetuning. Note that, the variables of augmentations applied to two frames keep the identical. We use the AdamW optimizer [28] with the one-cycle policy to train the model. The max learning rate is set to $3e^{-3}$. All experiments are performed using 8 NVIDIA Tesla V100.
Pre-training. The model is pre-trained for 12 epochs on the full training split of Waymo dataset, which contains 158,081 frames. The batch size is set to 4. Following GD-MAE [52], the masking ratio is set to 0.75, the number of predicted points per pillar is set as 16 and the number of target reconstruction points is set as $K^{GT}=64$.
Finetuning and inference. The model is finetuned for 30 epochs on different subsets of the Waymo training split. The number of epochs is held constant at 30, with a batch size of 2, while the data quantity varies. In addition to the data augmentations mentioned in the implementation details of Sec 4.1, following MV-JAR [48] and GD-MAE [52], a copy-n-paste augmentation [51] is also employed to slightly address the issue of class imbalance. During inference, each frame is selected as a current frame for a complete evaluation. The previous third frame, based on our ablation study, is selected as the previous frame.
4.2 Results on Waymo
In this section, the main results of the Waymo dataset are presented. We assess our approach after fine-tuning SiamWCA with varying amounts of labeled data. We explore two methods to obtain a specific percentage of the training set, but all models are evaluated on the same validation set.
$\triangleright$ Uniform sampling is commonly used to obtain a subset of LiDAR sequence [53, 52, 25]. Specifically, frames from all sequences are concatenated, from which the target frames are sampled with a fixed interval.
Data Amount Initialization Overall Vehicle Pedestrian Cyclist
L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH
5% Random 43.68 40.29 54.05 53.50 53.45 44.76 23.54 22.61
PointContrast [47] 45.32 41.30 52.12 51.61 53.68 43.22 30.16 29.09
ProposalContrast [53] 46.62 42.58 52.67 52.19 54.31 43.82 32.87 31.72
MV-JAR [48] 50.52 46.68 56.47 56.01 57.65 47.69 37.44 36.33
GD-MAE [$*{}^{*}$ 48.23 44.56 56.34 55.76 55.62 46.22 32.72 31.69
T-MAE (Ours) 50.20${}^{+6.52}$ 48.07${}^{+7.78}$ 56.71 56.18 60.19 55.71 33.69 32.33
10% Random 56.05 53.13 59.78 59.27 60.08 53.04 48.28 47.08
PointContrast [47] 53.69 49.94 54.76 54.30 59.75 50.12 46.57 45.39
ProposalContrast [53] 53.89 50.13 55.18 54.71 60.01 50.39 46.48 45.28
MV-JAR [48] 57.44 54.06 58.43 58.00 63.28 54.66 50.63 49.52
GD-MAE [$*{}^{*}$ 57.67 54.31 59.72 59.19 60.43 52.21 52.85 51.52
T-MAE (Ours) 59.14 ${}^{+3.09}$ 57.05 ${}^{+3.92}$ 60.29 59.79 63.61 59.24 53.52 52.13
20% Random 60.21 57.61 61.58 61.08 64.63 58.41 54.42 53.33
PointContrast [47] 59.35 55.78 58.64 58.18 64.39 55.43 55.02 53.73
ProposalContrast [53] 59.52 55.91 58.69 58.22 64.53 55.45 55.36 54.07
MV-JAR [48] 62.28 59.15 61.88 61.45 66.98 59.02 57.98 57.00
GD-MAE [$*{}^{*}$ 62.32 59.09 62.27 61.79 66.12 58.06 58.57 57.42
T-MAE (Ours) 63.48 ${}^{+3.27}$ 61.72 ${}^{+4.11}$ 62.93 62.47 68.26 64.53 59.24 58.16
20% Uniform Sampling Random 69.03 66.48 66.61 66.15 72.65 66.76 67.83 66.54
ProposalContrast [53] 66.67 64.20 65.22 64.80 66.40 60.49 68.48 67.38
BEV-MAE [26] 66.70 64.25 64.72 64.22 66.20 60.59 69.11 67.93
Occupancy-MAE [31] 65.86 63.23 64.05 63.53 65.78 59.62 67.76 66.53
MAELi-MAE [22] 65.60 63.00 64.22 63.70 65.93 59.79 66.66 65.52
GD-MAE [${}^{**}$ 70.24 67.14 67.67 67.22 73.18 65.50 69.87 68.71
AD-PT [55] 67.17 64.65 65.33 64.83 67.16 61.20 69.39 68.25
T-MAE (Ours) 70.61 ${}^{+1.58}$ 68.80 ${}^{+2.32}$ 66.85 66.39 74.85 71.08 70.14 68.94
100% Random 71.30 69.13 69.05 68.62 73.77 68.80 71.09 69.97
GCC-3D [25] 65.29 62.79 63.97 63.47 64.23 58.47 67.88 66.44
BEV-MAE [26] 66.92 64.45 64.78 64.29 66.25 60.53 69.73 68.52
PointContrast [47] 68.06 64.84 64.24 63.82 71.92 63.81 68.03 66.89
ProposalContrast [53] 68.17 65.01 64.42 64.00 71.94 63.94 68.16 67.10
MV-JAR [48] 69.16 66.20 65.52 65.12 72.77 65.28 69.19 68.20
GD-MAE [${}^{**}$ 70.62 67.64 68.72 68.29 72.84 65.47 70.30 69.16
T-MAE (Ours) 71.96 ${}^{+0.65}$ 70.22 ${}^{+1.09}$ 69.07 68.63 74.74 71.03 72.06 70.99
Table 1: Performance comparison with SSL methods on the Waymo validation set [39]. Random initialization denotes training frame scratch. 20% Uniform Sampling indicates sampling frames per five frames. $*{}^{*}$ indicates reproduced by us. ${}^{**}$ indicates from [52]. Results for other methods are taken from MV-JAR [48] or the survey [14]. Best results are highlighted as first, second, and third. Differences between T-MAE pre-training and random initialization are highlighted in red. T-MAE achieves significant improvements compared to random initialization and SOTA results compared to other SSL methods.
$\triangleright$ Data-efficient benchmark is proposed in MV-JAR [48] to solve the data diversity issue existing in the strategy of uniform sampling. More specifically, while trained on uniformly sampled frames, the models always converge to a similar performance if they are trained for adequate iterations no matter how many percentages of labeled data are used. As a result, fewer data do not lead to a shorter finetuning time. Therefore, to evaluate pre-trained representation more efficiently, we conduct experiments on the data-efficient benchmark, as presented in Tab. 1. Note that, for each data amount level, three non-overlap splits are provided MV-JAR [48]. The results for models finetuned by the second or the third split are presented in the supplementary material, as well as the average performance.
$\triangleright$ T-MAE achieves new SOTA at labelled data level of 10%, 20%, 20% (Uniform sampling) and 100%. Moreover, T-MAE achieves superior performance for the majority of class-specific metrics. To reduce performance variance on 5% and 10% splits, we also report in the supplementary material the results of T-MAE by finetuning it on the other two splits provided by [52]. Notably, thanks to the temporal modeling ability, T-MAE significantly outperforms other methods in terms of L2 mAPH for pedestrians. For instance, with 5% labeled data, T-MAE achieve better mAPH (i.e. 55.71) than any other method using 10% labeled data. The comparison with other SSL methods shows the efficacy of T-MAE. Nevertheless, MV-JAR [48] performs notably better for cyclists at 5% level and surpasses T-MAE in terms of L2 overall mAP.
$\triangleright$ The impact of the T-MAE pre-training becomes more pronounced when the labeled data is scarce. SiamWCA is a powerful backbone as, when finetuned with 100% of labeled data, it even performs better than other models that are initialized by different pre-training strategy (i.e. 69.13 v.s. 67.64). However, its performance drops significantly when finetuning data is limited (i.e. 40.29 v.s. 46.68 at 5% level). T-MAE consistently enhances model performance compared to random initialization, with more significant improvements observed when labeled data is limited.
4.3 Ablation study
In this section, we perform a series of ablation studies to assess the necessity of various components and justify the selection of hyperparameters. To cost-effectively experiment, the split 0 with 5% data of the data-efficient benchmark is used to finetune the model by default.
Is a delicate fusion module necessary? To compare with a model taking two frames as input, we modified the input of GD-MAE [52] by concatenating all points of two consecutive frames. We pre-trained, finetuned, and evaluated the modified GD-MAE with the same setting as T-MAE. As depicted in Table 2, directly merging two frames improves the metrics in terms of vehicles and pedestrians, which probably results from the density of target objects being doubled, as shown in Fig. 3. However, it has a negative impact on cyclists, resulting in a decrease in overall performance. The drop for cyclists might be attributed to the drift of estimated bounding boxes caused by the fast velocity of bicycles and their relatively small dimensions. In contrast, T-MAE significantly improves the performance thanks to the delicate pre-training strategy and fusion mechanism.
Method Frame Input L2 Overall Vehicle Pedestrian Cyclist
Previous Current mAP mAPH AP APH AP APH AP APH
GD-MAE - 0 48.23 44.56 56.34 55.76 55.62 46.22 32.72 31.69
GD-MAE - {-1, 0} 47.35 43.69 57.04 56.48 57.76 48.48 27.26 26.12
T-MAE (Ours) -1 0 49.45 46.78 56.56 56.02 57.96 51.94 33.82 32.37
Table 2: Two frames comparison. Two consecutive frames are merged and input into the GD-MAE [52] as an enhanced baseline.
Model Encoder Fusion L2 Overall Vehicle Pedestrian Cyclist
mAP mAPH AP APH AP APH AP APH
(a) Asymmetric WCA 47.26 44.78 54.25 53.75 56.66 50.98 30.88 29.60
(b) SimSiam WCA 45.58 42.05 53.37 52.84 53.80 44.78 29.58 28.53
(c) Siamese WCA$+$WSA 47.01 45.11 55.81 55.33 58.95 54.69 26.26 25.31
(d) Siamese ST WCA$+$WSA 46.29 43.64 54.13 53.64 55.73 49.24 29.02 28.04
(e) Ours Siamese WCA 49.45 46.78 56.56 56.02 57.96 51.94 33.82 32.37
Table 3: Ablation study on model architecture. The proposed SiamWCA demonstrates superior performance in terms of overall mAP and mAPH.
Encoder design and fusion mechanism. The proposed architecture consists of a Siamese encoder and a WCA module, which is denoted as (e) in this ablation study. In addition, several variants are compared: (a) Asymmetric encoder: The encoder for the previous frame is replaced with a more lightweight encoder, resulting in a reduction of the number of channels by half. Consequently, the two encoders no longer share weights. (b) We employ the SimSiam [7] approach, where one of the two encoders is disconnected from the computational graph, preventing it from being updated by gradient propagation. The encoder for the current frame is updated by gradients and then shares its weights with the detached encoder. (c) The design of the Siamese encoder stays unaltered, while a typical cross-self attention module is adopted in a window-based manner. These three variants are depicted in Fig. 5. (d) Instead of using GD-MAE [52] as our single-frame backbone, the standard (ST) MAE [18] is adopted to 3D space and serves as the Siamese encoder. The superiority of the proposed SiamWCA is shown in Tab. 3. Variant (c) outperforms SiamWCA in terms of pedestrians but at the expense of cyclists’ performance.
(a) Asym. Enc. (b) SimSiam Enc. (c) WCA + WSA
Figure 5: Three architecture variants of our SiamWCA backbone. (a) Asymmetric encoders: the encoder for the previous frame is replaced by a more lightweight encoder. (b) SimSiam-style [7] encoder: one encoder receives no gradient updates. (c) An additional windows-based self-attention is attached as the classic Transformer [43].
Previous frame L2 Overall Vehicle Pedestrian Cyclist
mAP mAPH AP APH AP APH AP APH
$[-5,-1]$ 49.86 47.61 56.70 56.16 59.47 54.65 33.41 32.03
$-1$ 49.45 46.78 56.56 56.02 57.96 51.94 33.82 32.37
$-3$ 50.20 48.07 56.71 56.18 60.19 55.71 33.69 32.33
$-5$ 49.87 47.69 56.73 56.20 59.91 55.26 32.98 31.61
Table 4: Temporal gap for inference. Different temporal gaps were tested for inference. $[-5,-1]$ indicates a random frame is sampled from the last five frames. Overall, a three-frame temporal gap worked best in our experiments.
Temporal interval for inference. This ablation study explores the impact of different temporal gaps during inference on model performance. We establish a baseline with a random interval between two frames. Specifically, for each frame, the previous frame is selected randomly from the preceding five frames. Furthermore, the temporal gap is fixed to $-1$, $-3$, and $-5$, respectively. If the number of available frames is less than the predefined threshold, the earliest available frame is selected. Tab. 4 illustrates that optimal performance is achieved with a fixed temporal gap of three. Specifically, the model excels in recognizing vehicles, pedestrians, and cyclists when the temporal gap is set to $-5$, $-3$, and $-1$, respectively. This observation highlights the varied impact of different temporal gaps on distinct object categories. Notably, while the performance remains relatively stable for vehicles and cyclists, there is a pronounced improvement in pedestrian recognition as the temporal gap transitions from $-1$ to $-3$. This highlights the considerable advantage of including more distant historical frames to achieve accurate pedestrian detection as well as his/her heading.
5 Conclusion
In this paper, we introduced the Temporal Mask AutoEncoders (T-MAE), a novel self-supervised paradigm for LiDAR point cloud pre-training. Building upon the single-frame MAE baseline, we incorporated historical frames into the representation using SiamWCA, with the proposed WCA module playing a pivotal role. This pre-training enabled the model to acquire robust representations and the ability to capture motion even with very limited labeled data. By considering temporally distant frames instead of consecutive ones, we achieved additional performance improvements. Our experiments on the Waymo Open dataset demonstrate the effectiveness of our approach by showing mostly improvements over state-of-the-art methods.
6 Acknowledgements
This work was financially supported by TomTom, the University of Amsterdam and the allowance of Top consortia for Knowledge and Innovation (TKIs) from the Netherlands Ministry of Economic Affairs and Climate Policy. Fatemeh Karimi Nejadasl was financed by the University of Amsterdam Data Science Centre.
Data Initialization Overall Vehicle Pedestrian Cyclist
amount L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH
5% Random 49.59 46.15 54.42 53.87 52.59 44.17 41.76 40.41
PointContrast [47] 48.97 44.91 52.35 51.85 52.49 41.95 42.07 40.91
ProposalContrast [53] 49.87 45.83 52.79 52.31 53.30 43.00 43.51 42.18
MV-JAR [48] 52.73 48.99 56.66 56.21 57.52 47.61 44.02 43.15
T-MAE (Ours) 52.91 ${}^{+3.32}$ 50.98 ${}^{+4.83}$ 55.60 55.05 57.98 54.18 45.14 43.70
10% Random 57.44 54.48 59.63 59.10 60.38 53.25 52.31 51.09
PointContrast [47] 55.22 51.31 55.62 55.15 59.25 49.17 50.81 49.60
ProposalContrast [53] 55.59 51.67 55.57 55.12 60.02 49.98 51.18 49.90
MV-JAR [48] 58.61 55.12 58.92 58.49 63.44 54.40 53.48 52.47
T-MAE (Ours) 59.18 ${}^{+1.74}$ 57.29 ${}^{+2.81}$ 59.66 59.15 63.90 59.87 53.97 52.84
Table 5: Performance comparison on the Waymo validation set [39]. Results for other methods are taken from MV-JAR [48]. All methods are finetuned with Subset 1.
Data Initialization Overall Vehicle Pedestrian Cyclist
amount L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH
5% Random 43.33 40.10 55.39 54.81 55.75 47.69 18.86 17.79
PointContrast [47] 44.48 40.55 51.87 51.37 55.36 45.03 26.22 25.24
ProposalContrast [53] 45.21 41.45 52.29 51.82 56.23 46.28 27.10 26.24
MV-JAR [48] 47.93 44.50 56.22 55.78 58.80 49.77 28.75 27.95
T-MAE (Ours) 46.89 ${}^{+3.56}$ 45.04 ${}^{+4.95}$ 56.32 55.81 60.90 56.65 23.45 22.67
10% Random 56.81 53.97 59.49 58.98 62.44 55.54 48.50 47.40
PointContrast [47] 54.80 51.02 55.41 54.95 60.56 50.86 48.44 47.24
ProposalContrast [53] 54.77 51.09 55.64 55.20 60.54 51.16 48.14 46.92
MV-JAR [48] 58.29 54.99 59.17 58.74 64.58 56.02 51.12 50.20
T-MAE (Ours) 58.15 ${}^{+1.34}$ 56.26 ${}^{+2.29}$ 60.08 59.56 66.02 61.82 48.34 47.41
Table 6: Performance comparison on the Waymo validation set [39]. Results for other methods are taken from MV-JAR [48]. All methods are finetuned with Subset 2.
Data Initialization Overall Vehicle Pedestrian Cyclist
amount L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH
5% Random 45.53 42.18 54.62 54.06 53.93 45.54 28.05 26.94
PointContrast [47] 46.26 42.25 52.11 51.61 53.84 43.40 32.82 31.74
ProposalContrast [53] 47.23 43.28 52.58 52.10 54.61 44.37 34.50 33.38
MV-JAR [48] 50.39 46.72 56.45 56.00 57.99 48.36 36.74 35.81
T-MAE (Ours) 50.00 ${}^{+4.46}$ 48.03 ${}^{+5.85}$ 56.21 55.68 59.69 55.51 34.09 32.90
10% Random 56.77 53.86 59.63 59.12 60.97 53.94 49.70 48.52
PointContrast [47] 54.57 50.75 55.26 54.80 59.85 50.05 48.61 47.41
ProposalContrast [53] 54.75 50.96 55.47 55.01 60.19 50.51 48.60 47.37
MV-JAR [48] 58.12 54.72 58.84 58.41 63.77 55.03 51.74 50.73
T-MAE (Ours) 58.82 ${}^{+2.06}$ 56.87 ${}^{+3.01}$ 60.01 59.50 64.51 60.31 51.94 50.79
Table 7: Average performance on the Waymo validation set [39], averaged across the models finetuned with Subset 0--2. Results for other methods are taken from MV-JAR [48]. T-MAE consistently enhances model performance with a notable margin compared to random initialization. Furthermore, T-MAE attains SOTA average performance across all metrics when finetuned with 10% data.
(a) Input (b) Attention scores for prior scan
Figure 6: Visualization of attention scores. (a) The input consists of two point clouds: the entire previous scan (red points) and the current scan (green points) that contains only points within the ground-truth bounding boxes. Note that, the blue bounding boxes in (a) serve solely for visualization purposes and do not function as input. (b) The pillar-wise attention scores are visualized. The attention scores are derived from the WCA module and mapped to a colormap ranging from black to white. The primary attention is placed on the target objects from the previous scan. This implies that the WCA is able to locate corresponding objects. Notably, WCA is capable of accurately trace to the source object, which is manually indicated by a skyblue box, even in cases where the vehicle is moving.
Method Data amount Epochs Number of input scans (×$\times 10^{3}$ per epoch) Total iterations (×$\times 10^{5}$) Number of GT used (×$\times 10^{3}$ per epoch) L2 mAPH (Overall) L2 mAPH (Pedestrian)
MV-JAR [48] 5% 72 7.9 5.70 7.9 46.68 47.69
10% 60 15.8 9.50 15.8 54.06 54.66
20% 48 31.6 15.19 31.6 59.15 59.02
Random/T-MAE 5% 30 7.9 2.37 4.0 40.29/48.07 44.76/55.71
10% 30 15.8 4.75 7.9 53.13/57.05 53.04/59.24
20% 30 31.6 9.50 15.8 57.61/61.72 58.41/64.53
Table 8: Quantitative details on comparison of finetuning iterations. Our scheme requires much less iterations for finetuning compared to MV-JAR [48].
Appendix A Additional Results on Waymo Subsets
The data-efficient benchmark, proposed by MV-JAR [48], contains three splits, each of which contains four tiers of data, namely 5%, 10%, 20% and 50%. In the context of data organization, each split corresponds to a subdivision where a lower level of data is a subset of a higher level. There is no overlap observed between different splits for the data corresponding to the 5% and 10% subsets. In the main paper, the majority of experiments were performed by finetuning the model with the split 0. To alleviate the variance caused by being finetuned with a small amount of data, we further present the performance obtained when the model is finetuned using the remaining two splits, as shown in Tab. 5 and Tab. 6. Furthermore, the average results are presented in Tab. 7. In conclusion, it can be observed that the model performance exhibits variation when finetuned with different subsets. However, T-MAE consistently demonstrates a noteworthy improvement in model performance compared to random initialization. Additionally, T-MAE attains SOTA average performance across all metrics when finetuned using 10% data.
Appendix B Visualization of cross-attention scores
To understand the knowledge acquired by the SiamWCA during pre-training, we employ attention visualization to identify critical regions of the previous scan. In particular, the entire previous scan and the target objects of current scan are input into the SiamWCA. More precisely, the input to the model for the current scan solely consists of points contained within the 3D ground-truth bounding boxes. The purpose of removing other points is to identify the specific areas of emphasis within the WCA module when the queries are solely target objects, e.g. vehicles and pedestrians. As illustrated in Fig. 6 (b), the attention is mainly attached to target objects, indicating that the WCA successfully detects and localizes these entities in the previous frame. Furthermore, the presence of vehicular action is perceptible even if the WCA module is only trained with unlabeled data, indicating the effectiveness of our T-MAE pre-training strategy.
Appendix C Comparison of training iterations.
MV-JAR [48] applies varying numbers of epochs during finetuning, depending on the data amount. Rather than varying numbers, we employ a predetermined number of epochs, leading to significantly less iterations for finetuning without performance drop. Detailed information is presented in Tab. 8. The total number of iterations is calculated by multiplying the number of input scans by the number of epochs. As depicted in Fig. 1 and Tab. 8, T-MAE achieve better performance than MV-JAR [48] while requiring a smaller number of finetuning iterations. In addition, our scheme requires just one GT annotation every two input frames. As a result, the total number of iterations would be even smaller if it is calculated by multiplying the number of GT by the number of epochs.

### v3

@@
 Abstract
-The scarcity of annotated data in outdoor point cloud segmentation poses a significant obstacle in harnessing the modeling capabilities of advanced networks like transformers.
-Consequently, scholars have been actively investigating efficacious self-supervised pre-training strategies, e.g.
-contrasting learning and reconstruction-based pretext tasks.
+The scarcity of annotated data in LiDAR point cloud understanding hinders effective representation learning.
+Consequently, scholars have been actively investigating efficacious self-supervised pre-training paradigms.
 Nevertheless, temporal information, which is inherent in the LiDAR point cloud sequence, is consistently disregarded.
 To better utilize this property, we propose an effective pre-training strategy, namely Temporal Masked Auto-Encoders (T-MAE), which takes as input temporally adjacent frames and learns temporal dependency.
-A SiamWCA backbone, containing a Siamese encoder and a window-based cross-attention (WCA) module, is established for the two-frame input.
-Taking into account that the motion of an ego-vehicle alters the illumination angles of the same instance, temporal modeling also serves as a robust and natural data augmentation, enhancing the comprehension of target objects.
-Moreover, instead of utilizing consecutive frames, it is more cost-effective and powerful by using distant historical frames.
+A SiamWCA backbone, containing a Siamese encoder and a windowed cross-attention (WCA) module, is established for the two-frame input.
+Considering that the movement of an ego-vehicle alters the view of the same instance, temporal modeling also serves as a robust and natural data augmentation, enhancing the comprehension of target objects.
 SiamWCA is a powerful architecture but heavily relies on annotated data.
-With our T-MAE pre-training strategy, we achieve the best performance on the Waymo dataset among self-supervised learning methods.
-Comprehensive experiments are conducted to validate all components of our proposal.
-Upon acceptance, the source code will be made accessible.
+Our T-MAE pre-training strategy alleviates its demand for annotated data.
+Comprehensive experiments demonstrate that T-MAE achieves the best performance on both Waymo and ONCE datasets among competitive self-supervised approaches.
+Codes will be released here.
+Abstract
+This supplementary provides more details and analysis of our method.
+The implementation details are illustrated in Sec.
+S1.
+Additional ablation results are provided in Sec.
+S2.
+Section S3 discusses the selection of temporal gap during inference.
+More quantitative comparisons with other methods on the Waymo dataset [49] are depicted in Sec.
+S4.
+The transferability of T-MAE is verified in Sec.
+S5 and a comparison with multi-frame non-SSL methods is provided in Sec.
+S6.
+We analyze what the WCA module learns from the T-MAE pre-training in Sec.
+S7.
+Sec.
+S8 provides details about the comparison of finetuning iterations.
+Eventually, more qualitative results and limitations are presented in Sec.
+S9 and Sec.
+S10.
+Keywords:
+Self-supervised learning LiDAR point cloud 3D detection
 1 Introduction
-Figure 1: T-MAE performance on Waymo dataset [39].
-Our T-MAE pre-training achieves substantially better results than random initialization.
-It also achieve better results than the SOTA SSL method MV-JAR [48] with significantly fewer iterations.
-Each point triplet corresponds to a specific fine-tuned model variant and shows the performance difference with two arrows.
-From left to right, the triplets correspond to a model fine-tuned with 5%, 10% and 20% data.
-In the domains of language and vision, researchers have observed that as models become more complex, the available amount of labeled data is often insufficient to adequately train huge models [11, 18] e.g.
-Vision Transformer (ViT) [12].
+[…]
+[…]
+Figure 1: T-MAE performance on Waymo [49].
+Left: Each point triplet shows the performance differences to three models finetuned with the same data.
+The triplets show finetuned models with 8K, 16K, 32K labeled frames (left to right).
+Our T-MAE pre-training outperforms both random initialization and the SOTA SSL method MV-JAR [60] with significantly fewer iterations.
+Right: T-MAE yields higher mAPH for pedestrians when finetuned with half the labeled data than MV-JAR.
+As deep neural networks become more complex, the available amount of labeled data is often insufficient to adequately train huge models [14, 24], e.g., Vision Transformer (ViT) [15].
 Consequently, there is a growing interest in exploring self-supervised learning (SSL) approaches as a potential solution to overcome this limitation.
@@
 The same challenge extends to the domain of point clouds, where annotations are more costly and time-consuming to obtain [58, 17].
-For instance, a mere 10% and 0.8% of sweeps are annotated in the nuScene [2] and ONCE [29] datasets.
-This challenge makes pre-training for point cloud understanding a non-trivial endeavor.
-Prior works mainly focus on synthetic and isolated objects [50, 56, 54, 35, 27] and indoor scene understanding [19, 24, 49, 36].
+For instance, a mere 10% and 0.8% of frames are annotated in the nuScene [4] and ONCE [37] datasets, respectively.
+This challenge makes pre-training for point cloud understanding a non-trivial endeavour.
+Prior works mainly focus on synthetic and isolated objects [62, 71, 69, 43, 34, 28, 72] and indoor scene understanding [25, 31, 61, 45, 35].
 Transferring these methods to outdoor LiDAR points is challenging due to their sparsity and dynamic environmental conditions.
-At present, most self-supervised methods used for understanding point clouds in autonomous driving rely on contrastive learning [47, 53, 32, 34, 33].
+At present, most of the self-supervised methods used for understanding point clouds in autonomous driving rely on contrastive learning [59, 67, 40, 42, 41].
 These approaches model the similarity and dissimilarity between entities, such as segments [41, 57] and/or points [59].
 In the wake of masked image modeling as a pretext task [24], efforts have also been devoted to the reconstruction of masked points [39, 53, 64, 60].
-The main idea is randomly masking points or voxels and urging the network to infer the coordinates of points [52] and/or voxels [48] or other properties, e.g.
-occupancy [31, 1] and curvature [41].
+The main idea is randomly masking points or voxels and urging the network to infer the coordinates of points [64] and/or voxels [60] or other properties, e.g., occupancy [39, 2] and curvature [53].
 Nevertheless, these methodologies often operate within the confines of a single-frame scenario, disregarding the fact that LiDAR data is typically acquired on a frame-by-frame basis.
-In other words, the valuable semantic information that is introduced by temporally adjacent frames is often not exploited.
+In other words, the valuable semantic information in temporally adjacent frames is barely exploited.
 Several methods attempt to leverage temporal information [27, 57, 41, 32] by incorporating multi-frame input during the self-supervised phase but their core concepts remain grounded in contrastive learning.
-Specifically, the scans captured at different times are treated as augmented samples of the same scene, without including temporal correspondence into the modeling procedure.
-Therefore, we propose a new self-supervised method, namely T-MAE, to exploit the accumulated observations.
-The key idea of T-MAE is to learn a powerful representation for point cloud scans and integrate features from the past for present use.
+Specifically, the point clouds captured at different times are treated as augmented samples of the same scene, without including temporal correspondence into the modeling procedure.
+Therefore, we propose a new self-supervised paradigm, namely T-MAE, to exploit the accumulated observations.
 During the pre-training stage, the current scan is voxelized with a high masking ratio, while the previous scan is fed entirely to the encoder.
-Then, the reconstruction pretext task is performed by incorporating three types of information: voxel embeddings of the past scan, visible voxel embeddings of the current scan, and the position of masked voxels.
-In this way, the WCA module learns to incorporate historical information into the current frame using unlabeled data.
-The proposed T-MAE pre-training strategy endows the network with both a powerful representation for sparse point clouds and the capacity to strengthen the present by learning from the past.
+Then, the pretext task is to reconstruct the current scan by incorporating voxel embeddings of the past scan, visible voxel embeddings of the current scan, and the position of masked voxels.
+This way, the proposed windowed cross-attention module learns to incorporate historical information into the current frame using unlabeled data.
+The T-MAE pre-training strategy endows the network with both a powerful representation for sparse point clouds and the capacity to strengthen the present by learning from the past.
 As shown in Fig.
 1, the proposed T-MAE achieves higher overall and pedestrian-specific mAPH than the randomly initialized baseline, namely the same model but trained from scratch.
-Moreover, T-MAE also outperforms state-of-the-art MV-JAR [48] with over 1.6×1.6$\times$1.6 × fewer training iterations.
-Our contributions are summarized as follows:
-  • We propose T-MAE, a novel and effective SSL approach for sparse point clouds, that learns temporal modeling in the process of reconstructing masked points.
-  • We design a SiamWCA backbone, containing a Siamese encoder and a window-based sparse cross-attention (WCA) module, to incorporate historical information.
-  • Rather than concatenating multiple consecutive point cloud scans, our approach demonstrates that incorporating more distant historical frames provides greater advantages.
-  • Our experiments demonstrate the efficacy of T-MAE by attaining impressive results on the Waymo Open dataset.
-T-MAE achieves 55.71 mAPH for pedestrians with only 5% of the labeled data and surpasses the SOTA SSL approaches by a substantial margin across all levels of fine-tuning data amount.
+Moreover, T-MAE also outperforms state-of-the-art MV-JAR [60] with over 1.6×1.6$\times$1.6 × to 2.4×2.4$\times$2.4 × fewer finetuning iterations.
+Our contributions are summarized as follows: 1) We propose T-MAE, a novel and effective SSL approach for representation learning of sparse point clouds, that learns temporal modeling in the process of reconstructing masked points.
+2) We design a SiamWCA backbone, containing a Siamese encoder and a windowed sparse cross-attention (WCA) module, to incorporate historical information.
+3) Our experiments demonstrate the efficacy of T-MAE by attaining substantial improvements on the Waymo and ONCE datasets.
+Notably, T-MAE with 5% labeled data outperforms the SOTA SSL approach MV-JAR in terms of mAPH for pedestrians, even if MV-JAR employs 10% labeled data.
 2 Related Work
-2.1 Static Self-supervised Learning
+Static Self-supervised Learning.
 SSL for point clouds is a burgeoning field due to the scarcity of annotations.
 In general, the pipeline of SSL consists of two phases.
-During the initial stage, namely pre-training with unlabeled data, the network is trained using a pretext task.
-Subsequently, in downstream tasks such as segmentation and detection, the pre-trained weights are loaded to the backbone, and the backbone is attached with task-specific heads.
-The resulting model is further fine-tuned using annotated data.
-The initial research efforts are primarily directed towards the exploration of contrastive learning.
+First, the network is trained using a pretext task with unlabeled data.
+Second, in downstream tasks such as segmentation and detection, the pre-trained weights are loaded to the backbone, and the backbone is attached with task-specific heads.
+The resulting model is finetuned with annotated data.
+The initial research efforts are primarily directed towards contrastive learning.
 The underlying hypothesis is that an image or 3D scene demonstrates feature equivalence even after undergoing different transformations [11, 7].
@@
 GCC-3D [32] exploit sequential information to obtain pseudo instances and then perform contrasting on the instance level.
-SegContrast [32] firstly obtains segments by an unsupervised clustering and then contrasting segments between two transformed views.
-These methods allow the network to learn equivalence with regard to geometric transformations.
-However, contrastive learning-based SSL approaches usually suffer from careful tuning of hyperparameters and complicated pre- or post-processing to find correspondences.
+SegContrast [40] firstly obtains segments by an unsupervised clustering and then contrasts segments between two transformed views.
+These methods allow the network to learn equivalence regarding geometric transformations.
+However, contrastive SSL approaches usually suffer from careful tuning of hyperparameters and complicated pre- or post-processing to find correspondences.
 The focus has shifted to reconstructing masked points, following the success of masked image modeling [24, 56].
@@
 OccupancyMAE [39] classifies voxels to whether they are occupied.
-Geo-MAE [41] infers occupancy as well as the normal and the curvature of each masked voxel.
+Geo-MAE [53] infers occupancy as well as the normal and curvature of each masked voxel.
 The GD-MAE’s pretext task is to reconstruct a fixed number of points for masked voxels [64].
@@
 In this paper, we employ the same masking strategy as GD-MAE [64] but expand the knowledge source to include both the current visible voxels and voxels from a previous frame as a reference.
-2.2 Temporal Self-supervised Learning
-There are many SSL methods designed to address temporal or spatiotemporal tasks.
-In video-related tasks, such as MAE-ST [15], a random patch masking strategy is employed over consecutive frames.
-The model is tasked with recovering these masked patches while considering information from neighboring frames.
+Temporal Self-supervised Learning.
+There are many SSL methods designed to address temporal or spatiotemporal tasks in the video domain.
+MAE-ST [18] employs a random patch masking strategy over consecutive frames.
+The model is tasked with recovering these masked patches while considering information from adjacent frames.
 VideoMAE [54] retains the same reconstruction pipeline as the pretext task but exploits a tube masking strategy.
-SiameseMAE [17] learns object-centered representations with the help of cross-attention layers and an asymmetrical masking technique on consecutive frames.
-There are also content-based masking strategies, e.g.
-motion-guided masking [20, 30].
+SiameseMAE [22] learns object-centred representations with the help of cross-attention layers and an asymmetrical masking technique on consecutive frames.
+There are also content-based masking strategies, e.g., motion-guided masking [26, 38].
 The primary concept underlying these methods is acquiring an understanding of temporal dependency across frames through the process of reconstructing patches with reference to consecutive frames.
-While this concept is effectively used in video understanding, it is rarely applied to the learning of sparse point cloud representations.
+While this concept is effectively used in video understanding, it is rarely applied to learn sparse point cloud representations.
 This is because point clouds are typically handled on a frame-by-frame basis rather than being regarded as a temporal sequence.
-For instance, when using temporally adjacent scans as input during self-supervised learning, both STSSL [45] and TARL [33] use HDBSCAN [3] to obtain segments and apply self-supervision by minimizing the feature distance between the same segments or points between neighbor frames.
+For instance, when using temporally adjacent scans as input during self-supervised learning, both STSSL [57] and TARL [41] use HDBSCAN [5] to obtain segments and apply self-supervision by minimizing the feature distance between the same segments or points of neighbour frames.
 While they achieve impressive results, two issues persist 1) The use of elaborate and time-consuming pre-processing methods to obtain segments.
 2) A primary focus on object consistency across frames rather than understanding object motion.
-Unlike other approaches, this paper aims to reduce the reliance on complex preprocessing techniques and emphasizes the significance of enabling the model to establish temporal correspondence through self-supervised learning from unlabeled data.
+Unlike other approaches, we aim to reduce the reliance on complex preprocessing techniques and focus on enabling the model to establish temporal correspondence through self-supervised learning from unlabeled data.
 3 Method
-Figure 2: Overview of the proposed T-MAE pre-training.
-Two frames are sampled from a sequence of point clouds according to temporal batch-based strategy.
-The prior scan $\mathcal{P}^{t+t_{1}}$ is voxelized and directly fed into the Siamese windowed cross-attention (SiamWCA) backbone.
-The current scan $\mathcal{P}^{t+t_{2}}$ undergoes a masking process before fed into the backbone.
-Furthermore, pillar-wise tokens are computed by a Siamese hierarchical encoder.
-The two-way gray arrow indicates weight sharing.
-The WCA module takes as input the full tokens of the previous frame and the partial observation of the current frame and creates enhanced tokens.
-In the diffusion process, masked tokens are updated by their neighbor tokens from WCA by a group of convolution layers.
-Subsequently, a reconstruction head recovers masked points.
-During fine-tuning, the full SiamWCA backbone is initialized with pre-trained weights obtained from the pre-training phase.
-In this section, previous works are briefly discussed in Sec.
-3.1.
-Then, the proposed T-MAE pre-training strategy is demonstrated in Sec.
-3.2.
-Eventually, the window-based sparse cross-attention (WCA) module is elaborated in Sec.
-3.3.
+We first briefly discuss important preliminaries in Sec.
+3.1 from previous works although they mainly focus on the single-frame setting.
+To incorporate historical frames, we build up our framework as in Fig.
+3.
+The key component of the framework, SiamWCA, is introduced in 3.2.
+Then, the proposed T-MAE pre-training strategy is described in Sec.
+3.3 and the corresponding windowed sparse cross-attention (WCA) module is elaborated in Sec.
+3.4.
 3.1 Preliminaries
@@
 These features can be treated as a pseudo-image in a bird’s eye view.
-On top of the pillar-based representation, SST [13] proposes a sparse regional self-attention (SRA) module to address a challenge in applying ViT to sparse LiDAR points.
-SST divides the pillars into non-overlapping windows with a fixed size and then applies self-attention to the non-empty pillars within the same window.
+On top of the pillar-based representation, SST [16] proposes a sparse regional self-attention (SRA) module to address challenges in applying ViT to sparse LiDAR points.
+SST divides the pillars into non-overlapping windows and then applies self-attention within each window.
 The receptive field of each pillar is expanded through a window shift operation.
-T-MAE and many recent works [52, 48, 40] are built up on the basic block SRA.
-Temporal batch.
-A sweep of a point cloud at time $t$ is denoted by $\mathcal{P}^{t}=\{(p^{t}_{k})\}_{k=1}^{K}$, where $p_{k}$ is a 3D point $p_{k}\in\mathbb{R}^{3}$.
-Moreover, a sequence of point clouds is an ordered set of $\mathcal{P}^{t}$, which is denoted by $\mathcal{P}=\{\mathcal{P}^{1},\mathcal{P}^{2},\ldots,\mathcal{P}^{T}\}$.
-During pre-training and fine-tuning, two scans are needed feed the Siamese encoder.
-Due to the high sampling frequency of the LiDAR system, two consecutive frames usually contain redundant duplicate information, which is verified in Sec.
-5.
-Conversely, when there is a substantial time gap between two frames, their overlap may be limited, rendering the historical information less useful.
-Therefore, we follow TARL [33] to introduce the concept of a temporal batch so that the interval of two frames is constrained to an appropriate range.
-Specifically, we sample a batch of consecutive frames $\mathcal{B}^{t}=\{\mathcal{P}^{t+1},\mathcal{P}^{t+2},\ldots,\mathcal{P}^{t+n}\}$ from a sequence $\mathcal{P}$.
-Then, two frames, $\mathcal{P}^{t+t_{1}}$ and $\mathcal{P}^{t+t_{2}}$ are sampled from the first and last one-third of the batch.
-More specifically, $0\leq t_{1}<\frac{n}{3}$ and $\frac{2n}{3}\leq t_{2}$.
-3.2 Temporal Masked Autoencoder Pre-training
-Inspired by the SiamMAE [17] used in video understanding, we develop Temporal Masked Auto-Encoders, namely T-MAE, for self-supervised learning on LiDAR points.
-As shown in Fig.
-2, the core idea is to reconstruct the present frame based on the historical and partial observations of the current frame.
-In this way, the network is compelled to learn a powerful sparse representation as well as the capacity to effectively model motion.
-After the pre-training, the weights of the SiamWCA backbone are retained for the downstream tasks.
-In the subsequent subsections, we provide a detailed elaboration of these steps.
-Single-frame baseline.
-To perform the reconstruction pretext task, we follow the concept of the state-of-the-art SSL method GD-MAE [52] which operates on a single frame.
-The pipeline encompasses several key stages, including voxelization, masking, encoding, decoding, and reconstruction.
-More specifically, the initial step voxelizes a given point cloud by two linear layers and a voxel-wise average pooling.
-The voxel representation can be considered a pillar-based representation, owing to its infinite height.
-Next, masking is applied in a pillar-wise manner.
-Subsequently, the remaining visible pillars serve as input for the sparse pyramid transformer to generate high-level tokens.
-Next, the sparse tokens are transformed into a dense tensor by means of zero-padding.
-Subsequently, a series of dense convolutional layers are utilized to dilate the tokens of unmasked voxels to vacant regions.
-$t$ ${t-3}\ \sim\ t$
-Refer to caption Refer to caption
-(a) A stationary vehicle.
-Refer to caption Refer to caption (b) A moving vehicle.
-Figure 3: Comparison between single and four frame concatenation.
+This window partition and shift make the SRA module achieve a good balance of efficiency and accuracy.
+T-MAE and many recent works [64, 60, 50] are built upon the basic SRA block.
+Reconstruction pretext task and single-frame baseline.
+The original MAE [24] randomly masks patches of an image and employs a ViT to reconstruct the image.
+To perform this reconstruction pretext task in the LiDAR domain, we follow the concept of the state-of-the-art SSL method GD-MAE [64] which operates on a single frame.
+The framework encompasses several key stages, i.e.
+voxelization, masking, encoding, dense feature recovery, and reconstruction.
+Through this pretext task, GD-MAE acquires valuable weights for the encoder, which are later utilized for downstream tasks.
+$t$ ${t-3}\ \sim\ t$ $t$ ${t-3}\ \sim\ t$
+Refer to caption Refer to caption Refer to caption Refer to caption
+(a) A stationary vehicle. (b) A moving vehicle.
+Figure 2: Comparison between single- and four-frame concatenation.
 While simple frame concatenation generally improves point density and detection rates, it can introduce spurious points in non-static scene parts that may degrade the detection performance.
@@
 The blue bounding boxes indicate the ground truth for the current frame.
-Figure 4: Window-based sparse cross-attention (WCA) module.
-Given the input tokens from both the previous and current frames, a joint region grouping is conducted to obtain a window partition.
-A sparse regional cross-attention (SRCA) is performed individually in each window to integrate the historical information to the middle tokens for the current frame.
-In other words, the tokens from two frames but with the same colors are attending to each other.
-For simplicity, the information flow is only illustrated for the green tokens.
-The grouping and cross-attention are performed once more with the shifted window partition.
-The red dot (Refer to caption) indicates the ego-vehicle driving towards the right.
-The box with diagonal stripes (Refer to caption) represents an object moving towards the left.
-The observed scene remains mostly static except for a pillar at bottom-left transitions from occupied to vacant due to occlusion and another pillar which does the opposite.
-Best viewed in color and high-resolution.
-SiamWCA.
+Analysis.
 Up to this point, a common approach has been to employ a single-frame reconstruction pretext task for LiDAR points.
@@
 However, as shown in Fig.
-3, while concatenation helps to identify static objects effectively, it introduces challenges for detectors when dealing with moving objects.
+2, while concatenation helps to identify static objects effectively, it introduces challenges for detectors when dealing with moving objects.
 We provide quantitative results in Tab.
-2.
-To address this problem, we suggest learning from past data in a feature space.
-Given two frames, the previous frame is transformed to the coordinate system of the later frame by their ego-poses.
-Pose information is usually readily available in relevant datasets [39, 2, 29] or easily obtained by means of GPS/IMU, odometry approaches [8], structure from motion (SfM) algorithms [37, 38], or SLAM systems [10].
+3.
+To address this problem, we suggest learning from past data in a latent space.
+3.2 Framework Overview
+Figure 3: Overview of our architecture and the proposed T-MAE pre-training.
+Two frames are sampled from a sequence of point clouds and are voxelized.
+During pre-training, the current frame $\mathcal{P}^{t_{2}}$ undergoes an additional masking process.
+Note that the dashed boxes indicate operations for pre-training phase only.
+Next, voxel-wise tokens are computed by a Siamese encoder.
+The two-way gray arrow indicates weight sharing.
+The WCA module takes as input the full tokens of the previous frame and the partial observation of the current frame and outputs enhanced tokens.
+The dense feature recovery places sparse tokens back to a dense feature map and convolves the map to fill empty locations.
+Subsequently, the feature map is either fed to a reconstruction head that recovers masked points, or to a detection head predicting bounding boxes.
+As illustrated in Fig.
+3, the proposed framework compromises voxelization, encoding, windowed cross-attention, dense feature recovery and two separate heads for pre-training and detection, respectively.
+This subsection elaborates on the key components for supervised learning, e.g., training from scratch and fine-tuning, whereas SSL-related components are introduced in Sec.
+3.3.
+Temporal batch-based sampling.
+Consider a sequence of point clouds as an ordered set of point clouds, denoted as $\mathbf{P}=\{\mathcal{P}^{1},\mathcal{P}^{2},\ldots,\mathcal{P}^{t},\ldots,\mathcal{P}^{T}\}$.
+In this ordered set, $\mathcal{P}^{t}=\{(p^{t}_{k})\}_{k=1}^{K}$ is a sweep of point cloud at time $t$, where $p_{k}$ is a 3D point $p_{k}\in\mathbb{R}^{3}$ and $K$ indicates the number of points.
+In our setting, two frames are needed to feed the network.
+Due to the high sampling frequency of the LiDAR system, two consecutive frames usually contain redundant duplicate information, which is verified in Sec.
+S3.
+Conversely, when there is a substantial time gap between two frames, their overlap may be limited, making the historical information less useful.
+Therefore, we follow TARL [41] to introduce the concept of a temporal batch so that the interval of two frames is constrained to an appropriate range.
+Specifically, we sample a batch of consecutive frames $\mathcal{B}^{t}=\{\mathcal{P}^{t+1},\mathcal{P}^{t+2},\ldots,\mathcal{P}^{t+n}\}$ from a sequence $\mathcal{P}$.
+Then, two frames, $\mathcal{P}^{t_{1}}$ and $\mathcal{P}^{t_{2}}$ are sampled from the first and last one-third of the batch, namely, $t_{1}\in\{t+1,t+2,\ldots,\left\lfloor t+\frac{n}{3}\right\rfloor\}$ and $t_{2}\in\{\left\lceil t+\frac{2n+1}{3}\right\rceil,\ldots,n-1,n\}$.
+Alignment and voxelization.
+Given two frames, the previous frame $\mathcal{P}^{t_{1}}$ is transformed to the coordinate system of the current frame $\mathcal{P}^{t_{2}}$ by their ego-poses.
+The pose information is available in most datasets [49, 4, 37] or easily obtained by means of GPS/IMU, odometry approaches [9], structure from motion (SfM) algorithms [46, 47], or SLAM systems [13].
+Next, each point cloud is divided into discrete voxels.
+Two linear layers map point coordinates to high-dimensional features and a voxel-wise representation is obtained via voxel-wise average pooling.
+The voxels can be regarded as pillars, owing to the infinite height.
+A Siamese encoder and windowed cross-attention (SiamWCA).
+A Siamese encoder [3] is a two-branch network where both branches share the same configurations and weights.
+In this work, it is utilized to encode the pillar-wise representations of both frames to sparse tokens.
+These tokens serve as input to the WCA module which facilitates the interaction between historical and current tokens.
+The SiamWCA is elaborated in Sec.
+3.4.
+Dense feature recovery and detection head.
+Once the current tokens are augmented with historical information, these sparse tokens are reverted to the x-y plane to form a dense feature map while vacant pillars are filled with zeros.
+Since LiDAR points only occur on object surfaces, the object centers typically locate at empty space, leading to inaccurate detection.
+We follow GD-MAE [64] to attach four dense convolutional layers, which spread the feature from occupied pillars to vacant regions.
+For the detection head, we adopt a center-based head and use the same target assignment strategy as CenterPoint [68].
+3.3 Temporal Masked Autoencoder Pre-training
+Up to this point, a two-frame framework for object detection has been set up.
+However, the transformer-based architecture is data-hungry.
+Inspired by the SiamMAE [22] used in video understanding, we develop Temporal Masked Auto-Encoders (T-MAE) for self-supervised learning on LiDAR points.
 As shown in Fig.
-2, the full observation of the previous frame is given to the network by passing it through voxelization and encoder.
-On the contrary, the later frame goes through voxelization, masking, and encoder.
-The encoders for the two frames share weights to ensure that the features are constrained in the same space.
-By utilizing the Siamese encoder, tokens are obtained for both frames.
-These tokens serve as input to the WCA module, which facilitates the interaction between historical and current tokens and is outlined in Sec.
-3.3.
-Once the current tokens are augmented with historical information, these sparse tokens are placed back into the x-y plane to form a dense tensor where vacant pillars are padded with zeros.
-Subsequently, a group of $3\times 3$ convolutional layers is utilized to process the tensor, which diffuses the features of visible pillars.
-Eventually, the obtained tokens of masked pillars are employed for point reconstruction.
-The Chamfer distance is computed as the loss function between the reconstructed points and the ground-truth points.
+3, the core idea is to reconstruct the present frame based on a full observation of the historical frame and a partial observation of the current frame.
+In this way, the network is compelled to learn a powerful sparse representation as well as the capacity to effectively model motion.
+After the pre-training, the weights of the SiamWCA backbone are retained for the downstream tasks.
+In the subsequent paragraphs, we elaborate on these steps.
+Masking.
+As shown in Fig.
+3, on top of the proposed SiamWCA backbone, an additional step, namely masking, is added between the voxelization and the encoder for the current frame.
+Specifically, masking is applied in a pillar-wise manner for a high ratio of the occupied pillars, e.g., 75%.
+The remaining pillars are subsequently input into the encoder.
+The encoders for the two frames share weights to ensure that the features are constrained in the same latent space.
+Due to masking, the number of tokens for the current frame decreases significantly, leading to much fewer valid windows and thus accelerating WCA during pre-training.
+The dense feature recovery performs exactly the same as in Sec.
+3.2 and outputs a dense feature map.
+Reconstruction head.
+Given the dense feature map, the reconstruction head retrieves the feature of masked pillars by their spatial location.
+With the pillar-wise features, the head reconstructs the relative coordinates of points, where the number of output points per pillar is set as $K^{O}$.
+Eventually, the Chamfer distance is computed as the loss function between the reconstructed points and the ground-truth points.
 Note that the number of points per pillar varies drastically.
 Thus, a fixed number of points $K^{GT}$ are randomly sampled as the target for reconstruction.
-In conclusion, SiamWCA, containing a Siamese encoder and a WCA module, computes the pillar-based representation for both frames and integrates the historical one to the present one for reconstruction.
-Unlike the single-frame baseline where only the encoder is reused for downstream tasks, the weights of SiamWCA are fully retained for downstream tasks.
-3.3 Window-based Sparse Cross-Attention Module
+In conclusion, T-MAE masks the voxelized tokens of the current frame and compels the network to reconstruct the current frame with the full observation of the historic frame as a reference, which encourages the WCA module to build up correspondence between frames.
+Note that, unlike the single-frame baseline where only the encoder is reused for downstream tasks, the weights of SiamWCA are fully retained for downstream tasks, which is proved to be effective as shown in Tab.
+S10.
+Therefore, the capability for building up correspondence is also retained.
+3.4 Siamese Encoder and Windowed Cross-Attention (SiamWCA)
+Figure 4: Windowed sparse cross-attention (WCA).
+Given the input tokens from both $\mathcal{P}^{t_{1}}$ and $\mathcal{P}^{t_{2}}$, a joint token grouping is performed to obtain a window partition.
+A sparse regional cross-attention (SRCA) is performed independently in each window to integrate the historical information to the middle tokens of the current frame.
+In other words, the tokens from two frames but with the same colors are attending to each other.
+For simplicity, the information flow is only depicted for the green tokens.
+After the second joint token grouping, the cross-attention are performed once more with the shifted window partition.
+The red dot (Refer to caption) indicates the ego-vehicle driving towards the right.
+The box with diagonal stripes (Refer to caption) represents an object, e.g., a vehicle, moving towards the left.
+Best viewed in color and high-resolution.
+Siamese encoder.
+Given the pillar-wise representations of a pair of frames, we explore an asymmetric network and a Siamese encoder for feature encoding.
+The asymmetric network consists of two branches with the same architecture but the branch for $\mathcal{P}^{t_{1}}$ is modified by reducing the number of channels by half, as depicted in Fig.
+6 (a).
+A Siamese encoder is a symmetric network with two subnetworks sharing weights.
+It is widely used to compute similarity in latent space for tracking [1, 23, 51, 21] and contrastive learning [10, 44, 12].
+For the Siamese encoder, we investigate two weight updating strategies, namely accumulation and SimSiam-style [10].
+Accumulation indicates the backward gradients of the two encoders are accumulated.
+SimSiam, standing for the simple Siamese pre-training strategy [10], indicates the encoder for $\mathcal{P}^{t_{1}}$ is detached from the computational graph, as depicted in Fig.
+6 (b).
+Thus the weight of this encoder is not updated by gradient propagation but by copying the weights from the encoder for $\mathcal{P}^{t_{2}}$.
+Windowed cross-attention (WCA).
 When provided with input tokens from two frames, there are several methods for interaction.
 One intuitive approach is to apply a standard self-attention layer to the concatenation of all tokens from both frames.
-This approach significantly increases GPU memory requirements.
-The vanilla Transformer block [43], which is combined with a cross-attention layer and a self-attention layer, can also be adopted.
-This combination, however, is demonstrated to be suboptimal as outlined in Sec.
-4.3.
-In due course, a single cross-attention layer is evaluated and exhibits superior performance compared to others.
-However, conventional global attention is not affordable in 3D space due to the expensive computational overhead imposed by the unavoidable high resolution.
-Therefore, the use of window-based attention is vital.
-The WCA module is illustrated in Fig.
+This approach significantly increases GPU memory requirements because it doubles the number of input tokens.
+The vanilla Transformer block [55], which consists of a cross-attention layer and a self-attention layer can also be adopted and a single cross-attention layer is feasible as well.
+However, all these conventional global attentions are not affordable in 3D space due to the expensive computational overhead imposed by the unavoidable high resolution.
+For instance, the input for our Siamese encoder can be $\mathbb{R}^{(468\times 468)\times 128}$, whereas it is $\mathbb{R}^{(14\times 14)\times 384}$ for a ViT-B/16 due to the patch embedding.
+Therefore, a window-based implementation is crucial for efficiency.
+In brief, the WCA module divides the 3D space into non-overlapping windows and then performs cross-attention within the windows.
+Figure 4 intuitively illustrates this process with a diagram that simulates a scene from a bird’s eye view.
+The key components are elaborated in the following paragraphs.
+$\triangleright$ Joint token grouping partitions the 3D space into non-overlapping windows and subsequently allocates the token to the corresponding window based on its spatial location.
+The previous frame has been aligned with the current frame and thus the window partition is unified for both frames, which allows the following attention mechanism to perform within each window.
+As shown in Fig.
+4, all tokens within the same physical window share the same colour, indicating that they are assigned to the same group for mutual attention.
+$\triangleright$ Sparse Regional Cross-Attention (SRCA) is essentially a cross-attention layer where the query comes from $\mathcal{P}^{t_{2}}$ and the key-value comes from $\mathcal{P}^{t_{1}}$.
+For clarity, we consider a single window as an example.
+Given two groups of tokens $\mathcal{F}^{t_{1}},\mathcal{F}^{t_{2}}$ from two frames and their corresponding spatial coordinates $\mathcal{I}^{t_{1}},\mathcal{I}^{t_{2}}$, the cross-attention is performed as follows:
+^$\hat{\mathcal{F}}^{t_{2}}$ $=\mathbf{MCA}\left(\mathcal{F}^{t_{2}}+\mathbf{PE}(\mathcal{I}^{t_{2}}),\mathcal{F}^{t_{1}}+\mathbf{PE}(\mathcal{I}^{t_{1}}),\mathcal{F}^{t_{1}}\right)$ (1)
+~$\widetilde{\mathcal{F}}^{t_{2}}$ $=\mathbf{LN}\left(\mathbf{MLP}\big{(}\mathbf{LN}(\hat{\mathcal{F}}^{t_{2}})\big{)}+\hat{\mathcal{F}}^{t_{2}}\right)+\mathcal{F}^{t_{2}}$ (2)
+where $\mathbf{MCA}(\mathcal{Q},\mathcal{K},\mathcal{V})$ indicates a classical Multi-head Cross-Attention, $\mathbf{PE}(\cdot)$ represents the absolute positional encoding function used in [6], and $\mathbf{LN}(\cdot)$ stands for Layer Normalization.
+Note that, if a window is empty in the previous frame, the tokens of this window in the current frame will remain unchanged, i.e.
+~$\widetilde{\mathcal{F}}^{t_{2}}=\mathcal{F}^{t_{2}}$.
+While the core concept of WCA is identical to cross-attention, this windowed implementation significantly scales down the computational complexity especially when working with sparse pillars.
+$\triangleright$ Repeated operations with window shift.
+Given the middle tokens, the window partition is shifted by half of the window size.
+Next, the tokens are re-grouped by the joint token grouping, as illustrated in “Shifted Window Partition” of Fig.
 4.
-The simulation involves an ego-vehicle moving to the right and another object moving in the opposite direction.
-The scene remains mostly static, but a pillar at bottom-left transitions from occupied to vacant due to occlusion.
-Joint Regional Grouping
-organizes the 3D spaces into non-overlapping windows.
-Hence, the following attention mechanism only performs with tokens from the same local region.
-The tokens from two frames are jointly considered so that the local windows across the two frames are corresponding.
-As shown in Fig.
-4, all tokens within the same physical window share the same color, indicating that they are assigned to the same group for mutual attention.
-Sparse Regional Cross-Attention
-(SRCA) is essentially a cross-attention layer where the query comes from the current frame and the key-value comes from the previous frame.
-For a group of tokens $\mathcal{F}^{t_{1}},\mathcal{F}^{t_{2}}$ and their corresponding spatial coordinates $\mathcal{I}^{t_{1}},\mathcal{I}^{t_{2}}$, the query $\mathcal{Q}$, key $\mathcal{K}$, and value $\mathcal{V}$ are calculated as follows
-$\mathcal{Q}$ $=\mathcal{F}^{t_{2}}+\mathbf{PE}(\mathcal{I}^{t_{2}})$ (1)
-$\mathcal{K}$ $=\mathcal{F}^{t_{1}}+\mathbf{PE}(\mathcal{I}^{t_{1}})$ (2)
-$\mathcal{V}$ $=\mathcal{F}^{t_{1}}$ (3)
-where $\mathbf{PE}(\cdot)$ represents the absolute positional encoding function used in [4]. SRCA follows conventional cross-attention and is defined as follows
-^$\hat{\mathcal{F}}^{t_{2}}$ $=\mathbf{MCA}(\mathcal{Q},\mathcal{K},\mathcal{V})$ (4)
-~$\widetilde{\mathcal{F}}^{t_{2}}$ $=\mathbf{LN}\left(\mathbf{MLP}\big{(}\mathbf{LN}(\hat{\mathcal{F}}^{t_{2}})\big{)}+\hat{\mathcal{F}}^{t_{2}}\right)$ (5)
-where $\mathbf{MCA}(\cdot,\cdot,\cdot)$ indicates the Multi-head Cross-Attention, and $\mathbf{LN}(\cdot)$ stands for Layer Normalization.
-While the core concept is similar to cross-attention, this implementation provides a more efficient solution when working with sparse pillars.
-Repeated Operations with Region Shift.
-The joint regional grouping and SRCA are performed once more after the regions are shifted.
-We follow the implementation of SST [13] to achieve region shift.
-Formally, suppose that the window size is $(l_{x},l_{y},l_{z})$, then the original windows are moved by $(l_{x}/2,l_{y}/2,l_{z}/2)$.
-Meanwhile, the tokens are re-grouped by the joint regional grouping, as illustrated in “Shifted Window Partition” of Fig.
-4.
-Note that, while performing the SRCA #2, the tokens ~$\widetilde{\mathcal{F}}^{t_{2}}$ of the current frame are updated by SRCA #1 while the tokens $\mathcal{F}^{t_{1}}$ remain unchanged.
+After that, the SRCA is performed once more.
+Note that, while performing the SRCA #2, the tokens ~$\widetilde{\mathcal{F}}^{t_{2}}$ of $\mathcal{P}^{t_{2}}$ are updated by SRCA #1 while the tokens $\mathcal{F}^{t_{1}}$ remain unchanged because a cross-attention operation does not update key $\mathcal{K}$ and value $\mathcal{V}$.
+With this window shift and SRCA #2, a token interacts with more tokens.
+For instance, the yellow token in the first window partition only attaches itself in the previous frame via SRCA #1 but interacts with more tokens in SRCA #2.
 4 Experiments
 4.1 Dataset and Implementation
-Waymo Open dataset [39] is the currently the largest public dataset with LiDAR points for autonomous driving.
-It comprises 798 sequences for training and 202 sequences for validation.
-Each sequence contains roughly 200 frames with pose information by which points can be transformed into the world coordinate system.
+Experiments are conducted on the following two datasets.
+Waymo Open dataset [49] is a large-scale autonomous driving dataset with LiDAR points.
 For 3D detection, an evaluation protocol is provided for calculating the average precision (AP) and the average precision weighted by heading (APH).
-Moreover, the evaluation includes two difficulty levels wherein bounding boxes containing over 5 points are regarded as Level 1 while Level 2 indicates all bounding boxes.
-We adopt mean AP and mean APH at Level 2 as main evaluation metrics.
+Moreover, the evaluation includes two difficulty levels wherein bounding boxes containing over 5 points are regarded as Level 1, while Level 2 indicates all bounding boxes.
+We adopt mean AP and mean APH at Level 2 as the main evaluation metrics.
+ONCE dataset [37] consists of 581 sequences of varying length.
+6, 4 and 10 sequences are selected for training, validation and test sets, respectively, and manually annotated.
+The remaining data is not annotated and adopted for pre-training in our experiments.
+The official evaluation metrics are AP calculated for each category and range.
 Implementation Details.
 We implement our approach based on the codebase of OpenPCDet [^1] https://github.com/open-mmlab/OpenPCDet.
-We keep the configuration of the sparse pyramid transformer consistent with its official implementation of GD-MAE [52].
-The window size of the WCA modules is set to $(8,8,1)$.
-The detection ranges for the Waymo dataset are restricted between $(-74.88m,-74.88m,-2m)$ and $(74.88m,74.88m,4m)$.
-The pillar size is $(0.32m,0.32m,6m)$.
-The size of the temporal batch is set to $n=6$ for both pre-training and finetuning.
+We adopt the sparse pyramid transformer as our encoder and keep the configuration consistent with its official implementation of GD-MAE [64].
 Three data augmentation techniques, i.e.
 random flipping, scaling, and rotation, are applied to both frames during pre-training and finetuning.
-Note that, the variables of augmentations applied to two frames keep the identical.
-We use the AdamW optimizer [28] with the one-cycle policy to train the model.
-The max learning rate is set to $3e^{-3}$.
-All experiments are performed using 8 NVIDIA Tesla V100.
-Pre-training.
-The model is pre-trained for 12 epochs on the full training split of Waymo dataset, which contains 158,081 frames.
-The batch size is set to 4.
-Following GD-MAE [52], the masking ratio is set to 0.75, the number of predicted points per pillar is set as 16 and the number of target reconstruction points is set as $K^{GT}=64$.
-Finetuning and inference.
-The model is finetuned for 30 epochs on different subsets of the Waymo training split.
-The number of epochs is held constant at 30, with a batch size of 2, while the data quantity varies.
-In addition to the data augmentations mentioned in the implementation details of Sec 4.1, following MV-JAR [48] and GD-MAE [52], a copy-n-paste augmentation [51] is also employed to slightly address the issue of class imbalance.
-During inference, each frame is selected as a current frame for a complete evaluation.
-The previous third frame, based on our ablation study, is selected as the previous frame.
-4.2 Results on Waymo
-In this section, the main results of the Waymo dataset are presented.
-We assess our approach after fine-tuning SiamWCA with varying amounts of labeled data.
-We explore two methods to obtain a specific percentage of the training set, but all models are evaluated on the same validation set.
-$\triangleright$ Uniform sampling is commonly used to obtain a subset of LiDAR sequence [53, 52, 25].
-Specifically, frames from all sequences are concatenated, from which the target frames are sampled with a fixed interval.
+In addition, following MV-JAR [60] and GD-MAE [64], a copy-n-paste augmentation [63] is also employed to slightly address the issue of class imbalance during finetuning.
+Note that identical data augmentations are applied to a pair of two frames.
+Further details are in the supplementary material.
+4.2 Main Results
+We present the comparison with SOTA methods on two datasets: For the ONCE dataset [37], we pre-train our SiamWCA with the raw_large split and fine-tune it with the annotated training set.
+For the Waymo dataset [49], we pre-train SiamWCA with the entire training set.
+Then, following MV-JAR [60], we finetune our approach with four portions of labeled data, namely 5%, 10%, 20% and 100%.
+Note that, a strong counterpart, GD-MAE [64], has not been evaluated in this setting and thus we re-train it for a more comprehensive comparison.
+[…]
 Data Amount Initialization Overall Vehicle Pedestrian Cyclist
-L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH
+mAP mAPH mAP mAPH mAP mAPH mAP mAPH
 5% Random 43.68 40.29 54.05 53.50 53.45 44.76 23.54 22.61
@@
 MV-JAR [60] 50.52 46.68 56.47 56.01 57.65 47.69 37.44 36.33
-GD-MAE [$*{}^{*}$ 48.23 44.56 56.34 55.76 55.62 46.22 32.72 31.69
-T-MAE (Ours) 50.20${}^{+6.52}$ 48.07${}^{+7.78}$ 56.71 56.18 60.19 55.71 33.69 32.33
+GD-MAE [64] 48.23 44.56 56.34 55.76 55.62 46.22 32.72 31.69
+T-MAE 50.89 47.22 57.06 56.05 58.95 52.62 36.64 32.99
+T-MAE (Ours) 51.47↑7.79 49.46↑9.17 57.13 56.63 59.69 55.28 37.61 36.48
 10% Random 56.05 53.13 59.78 59.27 60.08 53.04 48.28 47.08
@@
 MV-JAR [60] 57.44 54.06 58.43 58.00 63.28 54.66 50.63 49.52
-GD-MAE [$*{}^{*}$ 57.67 54.31 59.72 59.19 60.43 52.21 52.85 51.52
-T-MAE (Ours) 59.14 ${}^{+3.09}$ 57.05 ${}^{+3.92}$ 60.29 59.79 63.61 59.24 53.52 52.13
+GD-MAE [64] 57.67 54.31 59.72 59.19 60.43 52.21 52.85 51.52
+T-MAE 58.52 55.59 60.26 59.75 62.89 55.85 52.43 51.16
+T-MAE (Ours) 59.93↑3.88 57.99↑4.86 60.27 59.77 65.23 61.10 54.29 53.09
 20% Random 60.21 57.61 61.58 61.08 64.63 58.41 54.42 53.33
@@
 MV-JAR [60] 62.28 59.15 61.88 61.45 66.98 59.02 57.98 57.00
-GD-MAE [$*{}^{*}$ 62.32 59.09 62.27 61.79 66.12 58.06 58.57 57.42
-T-MAE (Ours) 63.48 ${}^{+3.27}$ 61.72 ${}^{+4.11}$ 62.93 62.47 68.26 64.53 59.24 58.16
-20% Uniform Sampling Random 69.03 66.48 66.61 66.15 72.65 66.76 67.83 66.54
-ProposalContrast [53] 66.67 64.20 65.22 64.80 66.40 60.49 68.48 67.38
-BEV-MAE [26] 66.70 64.25 64.72 64.22 66.20 60.59 69.11 67.93
-Occupancy-MAE [31] 65.86 63.23 64.05 63.53 65.78 59.62 67.76 66.53
-MAELi-MAE [22] 65.60 63.00 64.22 63.70 65.93 59.79 66.66 65.52
-GD-MAE [${}^{**}$ 70.24 67.14 67.67 67.22 73.18 65.50 69.87 68.71
-AD-PT [55] 67.17 64.65 65.33 64.83 67.16 61.20 69.39 68.25
-T-MAE (Ours) 70.61 ${}^{+1.58}$ 68.80 ${}^{+2.32}$ 66.85 66.39 74.85 71.08 70.14 68.94
+GD-MAE [64] 62.32 59.09 62.27 61.79 66.12 58.06 58.57 57.42
+T-MAE 62.37 60.17 62.19 61.72 67.18 62.18 57.74 56.59
+T-MAE (Ours) 63.52↑3.31 61.80↑4.19 63.10 62.59 68.23 64.66 59.23 58.15
 100% Random 71.30 69.13 69.05 68.62 73.77 68.80 71.09 69.97
@@
 MV-JAR [60] 69.16 66.20 65.52 65.12 72.77 65.28 69.19 68.20
-GD-MAE [${}^{**}$ 70.62 67.64 68.72 68.29 72.84 65.47 70.30 69.16
-T-MAE (Ours) 71.96 ${}^{+0.65}$ 70.22 ${}^{+1.09}$ 69.07 68.63 74.74 71.03 72.06 70.99
-Table 1: Performance comparison with SSL methods on the Waymo validation set [39].
-Random initialization denotes training frame scratch.
-20% Uniform Sampling indicates sampling frames per five frames.
-$*{}^{*}$ indicates reproduced by us.
-${}^{**}$ indicates from [52].
-Results for other methods are taken from MV-JAR [48] or the survey [14].
-Best results are highlighted as first, second, and third.
-Differences between T-MAE pre-training and random initialization are highlighted in red.
-T-MAE achieves significant improvements compared to random initialization and SOTA results compared to other SSL methods.
-$\triangleright$ Data-efficient benchmark is proposed in MV-JAR [48] to solve the data diversity issue existing in the strategy of uniform sampling.
-More specifically, while trained on uniformly sampled frames, the models always converge to a similar performance if they are trained for adequate iterations no matter how many percentages of labeled data are used.
-As a result, fewer data do not lead to a shorter finetuning time.
-Therefore, to evaluate pre-trained representation more efficiently, we conduct experiments on the data-efficient benchmark, as presented in Tab.
-1.
-Note that, for each data amount level, three non-overlap splits are provided MV-JAR [48].
-The results for models finetuned by the second or the third split are presented in the supplementary material, as well as the average performance.
-$\triangleright$ T-MAE achieves new SOTA at labelled data level of 10%, 20%, 20% (Uniform sampling) and 100%.
-Moreover, T-MAE achieves superior performance for the majority of class-specific metrics.
-To reduce performance variance on 5% and 10% splits, we also report in the supplementary material the results of T-MAE by finetuning it on the other two splits provided by [52].
+GD-MAE [64]∗∗ 70.62 67.64 68.72 68.29 72.84 65.47 70.30 69.16
+T-MAE 71.56 69.00 69.39 68.95 74.42 68.43 70.86 69.61
+T-MAE (Ours) 72.30↑1.00 70.52↑1.39 69.34 68.89 75.79 72.01 71.78 70.65
+The impact of the T-MAE pre-training.
+The bottom block in Tab.
+1 shows that the randomly initialized SiamWCA (denoted as Random) performs better than any other models that are initialized with a different pre-training strategy (i.e.
+69.13 v.s.
+67.64), suggesting that SiamWCA is a powerful backbone capable of learning temporal modeling when provided with sufficient annotated data.
+However, its performance drops significantly when finetuning data is limited (e.g., 40.29 v.s.
+46.68 at 5% level), indicating a strong demand for annotated data.
+As a comparison, the proposed T-MAE consistently enhances SiamWCA compared to random initialization.
+Moreover, as the labeled data shrinks from 100% to 5%, the impact of the T-MAE pre-training becomes more pronounced, namely increasing from 1.39 to 9.17, suggesting that the pre-training approach learns a powerful representation and alleviates the demand for annotated data.
+Figure 5: Qualitative results. We depict ground truth and predictions as boxes colored in red and green for two exemplary scenes from the Waymo dataset [49].
+Comparison with SOTA methods.
+We aim to leverage the temporal information between two adjacent frames, which is often overlooked by other methods.
+This absence makes it challenging to compare our method with others in the same setting.
+Therefore, we implement a test-time single-frame baseline (denoted as T-MAE) by replicating the same frame and inputting them into our pre-trained model during evaluation.
+Table 1 shows that T-MAE with identical frames outperforms SOTA counterparts.
+Moreover, T-MAE with adjacent frames achieves new SOTA at all levels in terms of overall and class-specific metrics.
 Notably, thanks to the temporal modeling ability, T-MAE significantly outperforms other methods in terms of L2 mAPH for pedestrians.
 For instance, with 5% labeled data, T-MAE achieve better mAPH (i.e.
-55.71) than any other method using 10% labeled data.
-The comparison with other SSL methods shows the efficacy of T-MAE.
-Nevertheless, MV-JAR [48] performs notably better for cyclists at 5% level and surpasses T-MAE in terms of L2 overall mAP.
-$\triangleright$ The impact of the T-MAE pre-training becomes more pronounced when the labeled data is scarce.
-SiamWCA is a powerful backbone as, when finetuned with 100% of labeled data, it even performs better than other models that are initialized by different pre-training strategy (i.e.
-69.13 v.s.
-67.64).
-However, its performance drops significantly when finetuning data is limited (i.e.
-40.29 v.s.
-46.68 at 5% level).
-T-MAE consistently enhances model performance compared to random initialization, with more significant improvements observed when labeled data is limited.
+55.28) than any other method using 10% labeled data.
+This metric indicates better direction detection for pedestrians, which could benefit downstream applications, e.g., pedestrian intention prediction.
+To reduce performance variance on 5% and 10% splits, we also report T-MAE performance in the Appendix (see Sec.
+S4) when it is finetuned with two more 5% and 10% splits and another 20% split.
+T-MAE consistently outperforms other methods, indicating the efficacy of T-MAE.
+Since there are no existing multi-frame SSL methods, we also compare our approach with robust non-SSL baselines that utilize multi-frame as inputs in the Appendix (see Tab.
+S9).
+To assess the generalization capabilities of our method, we also conducted experiments on the ONCE dataset [37].
+As shown in Tab.
+2, T-MAE outperforms other methods in most metrics, indicating its superiority.
+Moreover, the substantial improvement for pedestrians generalizes to this new dataset, indicating the dominance of T-MAE in pedestrian detection.
+Figure 5 shows two exemplary qualitative results from the Waymo dataset and more qualitative results are presented in the Appendix.
+Table 2: Performance comparisons on the validation split of the ONCE dataset [37].
+Pt.
+indicates the model is initialized with pre-trained weights.
+Results for other methods are taken from GD-MAE [64].
+Methods Pt. mAP Vehicle Pedestrian Cyclist
+Overall 0-30m 30-50m 50m-Inf Overall 0-30m 30-50m 50m-Inf Overall 0-30m 30-50m 50m-Inf
+PV-RCNN [48] 53.55 77.77 89.39 72.55 58.64 23.50 25.61 22.84 17.27 59.37 71.66 52.58 36.17
+IA-SSD [73] 57.43 70.30 83.01 62.84 47.01 39.82 47.45 32.75 18.99 62.17 73.78 56.31 39.53
+CenterPoint-Pillar [68] 59.07 74.10 85.23 69.22 53.14 40.94 48.43 34.72 20.09 62.17 73.70 56.05 40.19
+CenterPoint-Voxel [68] 60.05 66.79 80.10 59.55 43.39 49.90 56.24 42.61 26.27 63.45 74.28 57.94 41.48
+SECOND [63] 51.89 71.19 84.04 63.02 47.25 26.44 29.33 24.05 18.05 58.04 69.96 52.43 34.61
+w/ BYOL [20] 51.63 71.32 83.59 64.89 50.27 25.02 27.06 22.96 17.04 58.56 70.18 52.74 36.32
+w/ PointContrast [59] 53.59↑1.70 71.87 86.93 62.85 48.65 28.03 33.07 25.91 14.44 60.88 71.12 55.77 36.78
+w/ DeepCluster [52] 53.72↑1.83 72.89 83.52 67.09 50.38 30.32 34.76 26.43 18.33 57.94 69.18 52.42 34.36
+SPT [64] 62.62 75.64 87.21 70.10 53.21 45.92 54.78 37.84 22.56 66.30 78.12 60.52 42.05
+w/ GD-MAE [64] 64.92↑2.30 76.79 88.01 71.70 55.60 48.84 58.70 37.30 25.72 69.14 80.29 64.58 45.14
+SiamWCA (Ours) 63.71 76.47 87.63 71.59 55.16 47.27 57.57 36.99 21.79 67.40 78.39 62.78 43.90
+w/ T-MAE (Ours) 67.00↑3.29 78.35 88.45 73.05 57.16 52.57 62.66 44.18 25.29 70.09 81.14 65.33 46.48
 4.3 Ablation Study
-In this section, we perform a series of ablation studies to assess the necessity of various components and justify the selection of hyperparameters.
-To cost-effectively experiment, the split 0 with 5% data of the data-efficient benchmark is used to finetune the model by default.
+We perform ablation studies on the Waymo dataset [49] to justify design choices and hyperparameters.
+For cost-effective experiments, the split 0 with 5% data of the data-efficient benchmark [64] is used to finetune the model by default.
 Is a delicate fusion module necessary?
-To compare with a model taking two frames as input, we modified the input of GD-MAE [52] by concatenating all points of two consecutive frames.
-We pre-trained, finetuned, and evaluated the modified GD-MAE with the same setting as T-MAE.
-As depicted in Table 2, directly merging two frames improves the metrics in terms of vehicles and pedestrians, which probably results from the density of target objects being doubled, as shown in Fig.
-3.
-However, it has a negative impact on cyclists, resulting in a decrease in overall performance.
-The drop for cyclists might be attributed to the drift of estimated bounding boxes caused by the fast velocity of bicycles and their relatively small dimensions.
-In contrast, T-MAE significantly improves the performance thanks to the delicate pre-training strategy and fusion mechanism.
+Table 3: Two frames comparison. Two consecutive frames are merged and input into the GD-MAE [64] as an enhanced baseline.
 Method Frame Input L2 Overall Vehicle Pedestrian Cyclist
@@
 T-MAE (Ours) -1 0 49.45 46.78 56.56 56.02 57.96 51.94 33.82 32.37
-Table 2: Two frames comparison. Two consecutive frames are merged and input into the GD-MAE [52] as an enhanced baseline.
+To compare with a model taking two frames as input, we modified the input of GD-MAE [64] by concatenating all points of two consecutive frames.
+We pre-trained, finetuned, and evaluated the modified GD-MAE with the same setting as T-MAE.
+As depicted in Table 3, directly merging two frames improves the metrics in terms of vehicles and pedestrians, which probably results from the density of target objects being doubled, as shown in Fig.
+2.
+However, it has a negative impact on cyclists, resulting in a decrease in overall performance.
+The drop for cyclists might be attributed to the drift of estimated bounding boxes caused by the fast velocity of bicycles and their relatively small dimensions.
+In contrast, integrating a historical frame by our method consistently improves overall and class-specific metrics compared to the single-frame baseline.
+Table 4: Ablation study on model architecture. The proposed SiamWCA (e) demonstrates superior performance in terms of overall mAP and mAPH.
 Model Encoder Fusion L2 Overall Vehicle Pedestrian Cyclist
@@
 (c) Siamese WCA$+$WSA 47.01 45.11 55.81 55.33 58.95 54.69 26.26 25.31
-(d) Siamese ST WCA$+$WSA 46.29 43.64 54.13 53.64 55.73 49.24 29.02 28.04
+(d) Siamese WSA 43.82 40.90 53.05 52.54 54.08 48.18 24.33 21.97
 (e) Ours Siamese WCA 49.45 46.78 56.56 56.02 57.96 51.94 33.82 32.37
-Table 3: Ablation study on model architecture. The proposed SiamWCA demonstrates superior performance in terms of overall mAP and mAPH.
-Encoder design and fusion mechanism.
-The proposed architecture consists of a Siamese encoder and a WCA module, which is denoted as (e) in this ablation study.
-In addition, several variants are compared: (a) Asymmetric encoder: The encoder for the previous frame is replaced with a more lightweight encoder, resulting in a reduction of the number of channels by half.
+(a) Asym. Enc. (b) SimSiam Enc. (c) WCA + WSA (d) WSA
+Figure 6: Four architecture variants of our SiamWCA backbone.
+(a) Asymmetric encoders: the encoder for the previous frame is scaled down.
+(b) SimSiam-style [10] encoder: one encoder receives no gradient updates.
+(c) An additional windows-based self-attention is attached as the classic Transformer [55].
+(d) The fusion is implemented by a concat-WSA-split operation.
+Backbone design.
+The proposed architecture consists of a Siamese encoder and a WCA module.
+We ablate these components with alternatives.
+(a) Asymmetric encoder: it derives from a Siamese encoder but the encoder for $\mathcal{P}^{1}$ is scaled down.
 Consequently, the two encoders no longer share weights.
-(b) We employ the SimSiam [7] approach, where one of the two encoders is disconnected from the computational graph, preventing it from being updated by gradient propagation.
+(b) We employ the SimSiam [10] approach, where one of the two encoders is detached from the computational graph, preventing it from being updated by gradient propagation.
 The encoder for the current frame is updated by gradients and then shares its weights with the detached encoder.
 (c) The design of the Siamese encoder stays unaltered, while a typical cross-self attention module is adopted in a window-based manner.
-These three variants are depicted in Fig.
-5.
-(d) Instead of using GD-MAE [52] as our single-frame backbone, the standard (ST) MAE [18] is adopted to 3D space and serves as the Siamese encoder.
+(d) The WCA module in our SiamWCA is replaced with a windowed self-attention (WSA) module.
+However, since self-attention only needs one input, the tokens of both frames have to be concatenated.
+Then, the enhanced tokens of the current frame are split from the output of WSA module.
+These variants are depicted in Fig.
+6.
 The superiority of the proposed SiamWCA is shown in Tab.
-3.
-Variant (c) outperforms SiamWCA in terms of pedestrians but at the expense of cyclists’ performance.
-(a) Asym. Enc. (b) SimSiam Enc. (c) WCA + WSA
-Figure 5: Three architecture variants of our SiamWCA backbone.
-(a) Asymmetric encoders: the encoder for the previous frame is replaced by a more lightweight encoder.
-(b) SimSiam-style [7] encoder: one encoder receives no gradient updates.
-(c) An additional windows-based self-attention is attached as the classic Transformer [43].
-Previous frame L2 Overall Vehicle Pedestrian Cyclist
-mAP mAPH AP APH AP APH AP APH
-$[-5,-1]$ 49.86 47.61 56.70 56.16 59.47 54.65 33.41 32.03
-$-1$ 49.45 46.78 56.56 56.02 57.96 51.94 33.82 32.37
-$-3$ 50.20 48.07 56.71 56.18 60.19 55.71 33.69 32.33
-$-5$ 49.87 47.69 56.73 56.20 59.91 55.26 32.98 31.61
-Table 4: Temporal gap for inference.
-Different temporal gaps were tested for inference.
-$[-5,-1]$ indicates a random frame is sampled from the last five frames.
-Overall, a three-frame temporal gap worked best in our experiments.
+4.
+Variant (c) outperforms SiamWCA in terms of pedestrians but at the expense of both cyclists’ performance and extra parameters.
+Therefore, based on the comprehensive comparison, a Siamese encoder and a WCA fusion module are selected.
+Compatibility.
+To verify that the proposed SSL method T-MAE is independent of a specific encoder or detector, we replace the encoder SPT [64] with two other encoders, i.e.
+SST [16] and SpCNN [19], and the detector with a two-stage detector, Graph R-CNN [65].
+Table 5 shows that T-MAE constantly improves performance by a significant margin, showcasing its compatibility.
+Table 5: Ablation study on the encoder and the detector.
+The original encoder SPT [64] is replaced with SST [16] and SpCNN [19] and the original detector enterPoint [68] is replaced with Graph R-CNN [65].
+Due to space limits, only overall performance is presented.
+Data Amount Encoder Detector Random Init. T-MAE (Ours)
+mAP mAPH mAP mAPH
+5% SPT CenterPoint 43.68 40.29 51.47 49.46
+SST CenterPoint 44.26 41.19 51.59 49.24
+SpCNN CenterPoint 46.02 43.31 52.08 49.95
+SPT Graph R-CNN 51.18 47.27 56.92 54.70
+100% SPT CenterPoint 71.30 69.13 72.30 70.52
+SPT Graph R-CNN 72.76 70.04 75.16 73.50
 Temporal interval for inference.
-This ablation study explores the impact of different temporal gaps during inference on model performance.
-We establish a baseline with a random interval between two frames.
-Specifically, for each frame, the previous frame is selected randomly from the preceding five frames.
-Furthermore, the temporal gap is fixed to $-1$, $-3$, and $-5$, respectively.
-If the number of available frames is less than the predefined threshold, the earliest available frame is selected.
-Tab.
-4 illustrates that optimal performance is achieved with a fixed temporal gap of three.
-Specifically, the model excels in recognizing vehicles, pedestrians, and cyclists when the temporal gap is set to $-5$, $-3$, and $-1$, respectively.
-This observation highlights the varied impact of different temporal gaps on distinct object categories.
-Notably, while the performance remains relatively stable for vehicles and cyclists, there is a pronounced improvement in pedestrian recognition as the temporal gap transitions from $-1$ to $-3$.
-This highlights the considerable advantage of including more distant historical frames to achieve accurate pedestrian detection as well as his/her heading.
+The temporal interval between two frames should be constrained, as previously discussed in Sec.
+3.2.
+To investigate this matter, we conducted experiments on the Waymo Dataset [49], suggesting that a fixed interval of 0.3 seconds is better than using two consecutive frames.
+Additional information is available in Appendix Sec.
+S3.
 5 Conclusion
-In this paper, we introduced the Temporal Mask AutoEncoders (T-MAE), a novel self-supervised paradigm for LiDAR point cloud pre-training.
+We introduced Temporal Masked Autoencoders (T-MAE), a novel self-supervised paradigm for LiDAR point cloud pre-training.
 Building upon the single-frame MAE baseline, we incorporated historical frames into the representation using SiamWCA, with the proposed WCA module playing a pivotal role.
 This pre-training enabled the model to acquire robust representations and the ability to capture motion even with very limited labeled data.
-By considering temporally distant frames instead of consecutive ones, we achieved additional performance improvements.
-Our experiments on the Waymo Open dataset demonstrate the effectiveness of our approach by showing mostly improvements over state-of-the-art methods.
-6 Acknowledgements
+By constraining the temporal interval of two frames, we achieved additional performance improvement.
+Our experiments on the Waymo dataset and the ONCE dataset demonstrate the effectiveness of our approach by showing improvements over state-of-the-art methods.
+Acknowledgements
 This work was financially supported by TomTom, the University of Amsterdam and the allowance of Top consortia for Knowledge and Innovation (TKIs) from the Netherlands Ministry of Economic Affairs and Climate Policy.
+This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no.
+EINF-7940.
 Fatemeh Karimi Nejadasl was financed by the University of Amsterdam Data Science Centre.
-Data Initialization Overall Vehicle Pedestrian Cyclist
-amount L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH
-5% Random 49.59 46.15 54.42 53.87 52.59 44.17 41.76 40.41
+S1 Implementation Details
+We follow the training settings of GD-MAE [64].
+Some important configurations are listed in Tab.
+S1 and Tab.
+S2.
+We use the same masking ratio, the per-pillar number of predicted points $K^{O}$ and target reconstructed points $K^{GT}$ as GD-MAE because these parameters have negligible effects according to the ablation studies in GD-MAE [64].
+Table S1: Training details.
+Config Pre-training Finetuning
+optimizer AdamW [36]
+optimizer momentum $\beta_{1},\beta_{2}=0.9,0.99$
+weight decay 0.01
+max learning rate 0.003
+learning rate scheduler a cyclic learning rate
+with cosine annealing
+batch size (Waymo [49]) 4 3
+batch size (ONCE [37]) 8 6
+epoch 12 30
+masking ratio 0.75 -
+# predicted points $K^{O}$ 16 -
+# target reconstructed points $K^{GT}$ 64 -
+Table S2: Dataset-specific details. Finetuning time indicates the duration of finetuning using the entire training set.
+Config Waymo [49] ONCE [37]
+window size $(8,8,1)$
+detection range - x-axis (m) $(-74.88,74.88)$
+detection range - y-axis (m) $(-74.88,74.88)$
+detection range - z-axis (m) $(-2,4)$ $(-3,5)$
+pillar size (m) $(0.32,0.32,6)$ $(0.32,0.32,8)$
+temporal batch 6 3
+GPUs 8 $\times$ Tesla V100 4 $\times$ A100 (40GB)
+Pre-training time (GPU day) 8 $\times$ 4 4 $\times$ 4.5
+Finetuning time (GPU day) 8 $\times$ 5 4 $\times$ 0.25
+Table S3: Ablation experiments on the Waymo dataset [49].
+Ablation Target Setting L2 Overall
+mAP mAPH
+Two-frame alignment baseline 41.42 37.56
+w/ alignment 44.05↑2.63 41.28↑3.72
+Data augmentation w/ copy-n-paste [63] 46.76↑5.34 43.93↑6.37
+Length of temporal batch 6$\rightarrow$3 45.88↑4.46 41.96↑4.40
+6 46.76↑5.34 43.93↑6.37
+6$\rightarrow$12 44.96↑3.54 41.92↑4.36
+S2 Additional Ablation Study
+Additional results on the Waymo dataset [49] are provided to justify design choices and hyperparameter values.
+Note that, in this section, we use 20% of the training set to pre-train our model and use 5% of the training set to fine-tune it for cost-effective experiments.
+Two-frame Alignment.
+We start with using SiamWCA as a baseline and investigate whether it is necessary to align the previous frame to the coordinate system of the current frame.
+As shown in Tab.
+S3, aligning two frames significantly improves the metrics.
+Data Augmentation.
+Based on the baseline with two-frame alignment, we further investigate the effect of a data augmentation technique, namely copy-n-paste [63].
+Specifically, it first generates a ground truth instance database during dataset pre-processing.
+Then, during finetuning, several ground truth instances are randomly placed into the scene.
+This augmentation boosts the detection rates.
+Note that only ground truths in the 5% split are included in the database, which avoids ground truth leakage.
+Length of Temporal Batch.
+This hyperparameter is set as 6 for the previous experiments in Tab.
+S3.
+In this ablation, we pre-train and fine-tune the model with different lengths of the temporal batch.
+As shown in Tab.
+S3, setting the length of the temporal batch as 6 achieves the optimal performance.
+S3 Temporal interval for inference.
+Figure S1: Temporal interval for inference.
+Different temporal intervals were tested for inference.
+$[0.1,0.5]$ indicates a random interval between $0.1$ and $0.5$ seconds.
+A fixed interval ranging from 0.1 to 0.9 seconds was also tested.
+Overall, a fixed interval of 0.3 seconds works best in our experiments.
+This ablation study explores the impact of different temporal intervals on the model performance during inference.
+We establish a baseline with a random interval ranging from 0.1 to 0.5 seconds.
+In other words, $\mathcal{P}^{t_{1}}$ is selected randomly from the last five frame of $\mathcal{P}^{t_{2}}$.
+As a comparison, the temporal interval is fixed to values ranging from $0.1$ to $0.9$ seconds.
+If there is no satisfactory frame available, the earliest available frame is selected as $\mathcal{P}^{t_{1}}$.
+As shown in Fig.
+S1, The AP and APH for pedestrians achieve relatively high values when the interval is 0.3 and 0.5 seconds, while those for cyclists hit optimal when the interval is 0.1 seconds.
+The different optimal intervals could be attributed to the different speeds of the two categories.
+Pedestrians can barely move within 0.1 seconds, leading to historical information less useful.
+On the contrary, cyclists move relatively faster and thus the same bicycle of two frames is easy to be recognized as two instances if the temporal gap is large.
+Fig.
+S1 also illustrates that optimal overall performance is achieved when the interval is 0.3 seconds.
+Therefore, the previous third frame, which corresponds to a temporal interval of 0.3 seconds, is selected as $\mathcal{P}^{t_{1}}$ for the Waymo dataset.
+S4 Additional Results on Waymo dataset
+Table S4: Performance comparison on the Waymo validation set [49].
+All methods are fine-tuned with 20% uniformly sampled frames, namely sampling one frame per five frames.
+Random initialization denotes training from scratch.
+Differences between T-MAE pre-training and random initialization are highlighted in red.
+∗∗ indicates results from [64].
+Other results are from AD-PT [70] or the survey [17].
+Best results are highlighted as first, second, and third.
+Data Amount Initialization Overall Vehicle Pedestrian Cyclist
+mAP mAPH mAP mAPH mAP mAPH mAP mAPH
+20% Uniform Sampling Random 69.03 66.48 66.61 66.15 72.65 66.76 67.83 66.54
+ProposalContrast [67] 66.67 64.20 65.22 64.80 66.40 60.49 68.48 67.38
+BEV-MAE [33] 66.70 64.25 64.72 64.22 66.20 60.59 69.11 67.93
+Occupancy-MAE [39] 65.86 63.23 64.05 63.53 65.78 59.62 67.76 66.53
+MAELi-MAE [29] 65.60 63.00 64.22 63.70 65.93 59.79 66.66 65.52
+GD-MAE [64]∗∗ 70.24 67.14 67.67 67.22 73.18 65.50 69.87 68.71
+AD-PT [70] 67.17 64.65 65.33 64.83 67.16 61.20 69.39 68.25
+T-MAE (Ours) 70.92↑1.89 69.11↑2.63 68.07 67.61 74.38 70.56 70.32 69.15
+Table S5: Performance on the Waymo validation set [49]. Results for other methods are taken from MV-JAR [60]. All methods are finetuned with Subset 1.
+Finetuning Initialization Overall Vehicle Pedestrian Cyclist
+split mAP mAPH mAP mAPH mAP mAPH mAP mAPH
+5% S1 Random 49.59 46.15 54.42 53.87 52.59 44.17 41.76 40.41
 PointContrast [59] 48.97 44.91 52.35 51.85 52.49 41.95 42.07 40.91
@@
 MV-JAR [60] 52.73 48.99 56.66 56.21 57.52 47.61 44.02 43.15
-T-MAE (Ours) 52.91 ${}^{+3.32}$ 50.98 ${}^{+4.83}$ 55.60 55.05 57.98 54.18 45.14 43.70
-10% Random 57.44 54.48 59.63 59.10 60.38 53.25 52.31 51.09
+T-MAE (Ours) 53.44↑3.85 51.58↑5.43 56.18 55.65 57.92 54.17 46.22 44.90
+10% S1 Random 57.44 54.48 59.63 59.10 60.38 53.25 52.31 51.09
 PointContrast [59] 55.22 51.31 55.62 55.15 59.25 49.17 50.81 49.60
@@
 MV-JAR [60] 58.61 55.12 58.92 58.49 63.44 54.40 53.48 52.47
-T-MAE (Ours) 59.18 ${}^{+1.74}$ 57.29 ${}^{+2.81}$ 59.66 59.15 63.90 59.87 53.97 52.84
-Table 5: Performance comparison on the Waymo validation set [39]. Results for other methods are taken from MV-JAR [48]. All methods are finetuned with Subset 1.
-Data Initialization Overall Vehicle Pedestrian Cyclist
-amount L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH
-5% Random 43.33 40.10 55.39 54.81 55.75 47.69 18.86 17.79
+T-MAE (Ours) 59.24↑1.80 57.26↑2.78 59.71 59.19 64.44 60.29 53.58 52.31
+Table S6: Performance on the Waymo validation set [49]. Results for other methods are taken from MV-JAR [60]. All methods are finetuned with Subset 2.
+Finetuning Initialization Overall Vehicle Pedestrian Cyclist
+split mAP mAPH mAP mAPH mAP mAPH mAP mAPH
+5% S2 Random 44.25 41.48 53.04 52.53 55.16 49.42 24.56 22.48
 PointContrast [59] 44.48 40.55 51.87 51.37 55.36 45.03 26.22 25.24
@@
 MV-JAR [60] 47.93 44.50 56.22 55.78 58.80 49.77 28.75 27.95
-T-MAE (Ours) 46.89 ${}^{+3.56}$ 45.04 ${}^{+4.95}$ 56.32 55.81 60.90 56.65 23.45 22.67
-10% Random 56.81 53.97 59.49 58.98 62.44 55.54 48.50 47.40
+T-MAE (Ours) 49.01↑4.75 46.21↑4.74 56.50 55.95 60.70 54.70 29.82 27.98
+10% S2 Random 56.81 53.97 59.49 58.98 62.44 55.54 48.50 47.40
 PointContrast [59] 54.80 51.02 55.41 54.95 60.56 50.86 48.44 47.24
@@
 MV-JAR [60] 58.29 54.99 59.17 58.74 64.58 56.02 51.12 50.20
-T-MAE (Ours) 58.15 ${}^{+1.34}$ 56.26 ${}^{+2.29}$ 60.08 59.56 66.02 61.82 48.34 47.41
-Table 6: Performance comparison on the Waymo validation set [39]. Results for other methods are taken from MV-JAR [48]. All methods are finetuned with Subset 2.
-Data Initialization Overall Vehicle Pedestrian Cyclist
-amount L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH
-5% Random 45.53 42.18 54.62 54.06 53.93 45.54 28.05 26.94
+T-MAE (Ours) 58.64↑1.83 56.71↑2.73 60.20 59.70 66.10 61.79 49.62 48.63
+Table S7: Average performance on the Waymo validation set [49], averaged across the models finetuned with Subset 0$\sim$2.
+Results for other methods are taken from MV-JAR [60].
+T-MAE consistently enhances model performance with a notable margin compared to random initialization.
+Finetuing Initialization Overall Vehicle Pedestrian Cyclist
+split mAP mAPH mAP mAPH mAP mAPH mAP mAPH
+[…]
 PointContrast [59] 46.26 42.25 52.11 51.61 53.84 43.40 32.82 31.74
@@
 MV-JAR [60] 50.39 46.72 56.45 56.00 57.99 48.36 36.74 35.81
-T-MAE (Ours) 50.00 ${}^{+4.46}$ 48.03 ${}^{+5.85}$ 56.21 55.68 59.69 55.51 34.09 32.90
-10% Random 56.77 53.86 59.63 59.12 60.97 53.94 49.70 48.52
+T-MAE (Ours) 51.31↑5.47 49.08↑6.44 56.60 56.08 59.44 54.72 37.88 36.46
+[…]
 PointContrast [59] 54.57 50.75 55.26 54.80 59.85 50.05 48.61 47.41
@@
 MV-JAR [60] 58.12 54.72 58.84 58.41 63.77 55.03 51.74 50.73
-T-MAE (Ours) 58.82 ${}^{+2.06}$ 56.87 ${}^{+3.01}$ 60.01 59.50 64.51 60.31 51.94 50.79
-Table 7: Average performance on the Waymo validation set [39], averaged across the models finetuned with Subset 0--2.
-Results for other methods are taken from MV-JAR [48].
-T-MAE consistently enhances model performance with a notable margin compared to random initialization.
-Furthermore, T-MAE attains SOTA average performance across all metrics when finetuned with 10% data.
-(a) Input (b) Attention scores for prior scan
-Figure 6: Visualization of attention scores.
-(a) The input consists of two point clouds: the entire previous scan (red points) and the current scan (green points) that contains only points within the ground-truth bounding boxes.
+T-MAE (Ours) 59.27↑2.50 57.32↑3.46 60.06 59.55 65.26 61.06 52.50 51.34
+We explore two methods to obtain a subset of the training set for finetuning, but all models are evaluated on the same validation set.
+Uniform Sampling is commonly used [67, 64, 32].
+Specifically, frames from all sequences are concatenated, from which the target frames are sampled with a fixed interval.
+Data-efficient Benchmark [60] selects a certain number of sequences as a training subset instead of uniform sampling frames, which aims to solve the data diversity issue existing in uniform sampling.
+More specifically, while finetuned on uniformly sampled frames, models always converge to a similar performance if they are finetuned for adequate iterations no matter how many percentages of labelled data are used.
+As a result, fewer data do not lead to a shorter finetuning time.
+The data-efficient benchmark enables models to converge with much fewer iterations.
+Therefore, to evaluate pre-trained representation more efficiently, we conduct experiments on the data-efficient benchmark, as presented in Tab.
+1.
+Results Analysis.
+Under the uniform sampling setting, we finetuned the T-MAE pre-trained model with 20% uniformly sampled frames.
+As shown in Tab.
+S4, T-MAE outperforms other methods, which aligns with the statement made in the main paper.
+Under the data-efficient setting, the 5% and 10% splits contain limited samples, which may lead to performance variance.
+Therefore, except for the split 0 used for Tab.
+1, we also compare models in Tab.
+S5 and Tab.
+S6 when they are finetuned with split 1 and split 2, both of which are provided by MV-JAR [60] as well.
+Furthermore, the average results are presented in Tab.
+S7.
+In conclusion, it can be observed that the model performance exhibits variation when finetuned with different subsets.
+However, T-MAE pre-training consistently improves model performance compared to random initialization.
+T-MAE pre-training surpasses other SSL methods in terms of most class-specific metrics and all overall metrics.
+Notably, T-MAE outperforms other methods in terms of all metrics while the results are averaged (see Tab.
+S7).
+S5 Transferring performance.
+To assess the transferability of T-MAE pre-training, we pre-train the model on the ONCE dataset [37] and then fine-tune it on the Waymo dataset [49].
+Table S8 demonstrates that T-MAE offers robust generalizability and transferability.
+Table S8: Transferability comparison.
+Initialization Data Overall Vehicle Pedestrian Cyclist
+Random - 66.48 66.15 66.76 66.54
+GD-MAE [64] ONCE 66.61 67.18 64.82 67.83
+T-MAE (Ours) ONCE 67.86 68.06 67.25 68.26
+S6 Multi-frame comparison with non-SSL methods.
+Since there are no existing multi-frame SSL methods, we compare our approach with robust non-SSL baselines that utilize multi-frame as inputs.
+Table S9 demonstrates our method surpasses other supervised methods while requiring only two frames, highlighting the effectiveness of our method.
+Table S9: Multi-frame comparison with non-SSL methods. “-” indicates not available.
+Method Frames L2 Overall Vehicle Pedestrian Cyclist
+mAP mAPH mAP mAPH mAP mAPH mAP mAPH
+3D-MAN [66] 16 - - 67.6 67.1 62.6 59.0 - -
+CenterPoint [68] 4 70.8 69.4 69.1 68.6 71.7 68.6 71.6 70.9
+SST [16] 3 - - 68.5 68.1 75.1 70.9 - -
+T-MAE (Ours) 2 72.3 70.5 69.4 68.9 75.8 72.0 71.8 70.7
+(a) Input (b) Attention scores for prior frame
+Figure S2: Visualization of attention scores.
+(a) The input consists of two point clouds: the entire previous frame (red points) and the current frame (green points) that contains only points within the ground-truth bounding boxes.
 Note that, the blue bounding boxes in (a) serve solely for visualization purposes and do not function as input.
@@
 The attention scores are derived from the WCA module and mapped to a colormap ranging from black to white.
-The primary attention is placed on the target objects from the previous scan.
+The primary attention is placed on the target objects from the previous frame.
 This implies that the WCA is able to locate corresponding objects.
 Notably, WCA is capable of accurately trace to the source object, which is manually indicated by a skyblue box, even in cases where the vehicle is moving.
+S7 Attention Learned by T-MAE Pre-training
+In the main paper, T-MAE pre-trained weights are loaded to both the Siamese encoder and WCA module.
+We conduct an ablation study where only the Siamese encoder is initialized by the pre-trained weights and the WCA module is randomly intialized.
+As shown in Tab.
+S10, initializing WCA with T-MAE pre-trained weights significantly improves mAPH for pedestrians, which indicates a way better direction detection for pedestrians.
+It also boosts metrics for cyclists with a big margin.
+Table S10: Ablation on WCA initialization. SE stands for the Siamese Encoder.
+Initialization Pre-trained Overall Vehicle Pedestrian Cyclist
+SE WCA mAP mAPH mAP mAPH mAP mAPH mAP mAPH
+Random 43.68 40.29 54.05 53.50 53.45 44.76 23.54 22.61
+Partially random 48.19↑4.51 45.16↑4.87 55.91 55.38 56.29 48.74 32.38 31.36
+T-MAE (ours) 51.47↑7.79 49.46↑9.17 57.13 56.63 59.69 55.28 37.61 36.48
+To further understand the knowledge acquired by the WCA module during pre-training, we employ attention visualization to identify critical regions of the previous frame.
+In particular, the entire previous frame and the target objects of the current frame are input into the SiamWCA that loads T-MAE pre-trained weights.
+More precisely, the input to the model for the current frame solely consists of points contained within the 3D ground-truth bounding boxes.
+The purpose of removing other points is to identify the specific areas of emphasis within the WCA module when the queries are solely target objects, e.g.
+vehicles, pedestrians, and cyclists.
+As illustrated in Fig.
+S2 (b), the attention is mainly attached to target objects, indicating that the WCA successfully detects and localizes these entities in the previous frame.
+Furthermore, the presence of vehicular action is perceptible even if the WCA module is only trained with unlabeled data, indicating the effectiveness of our T-MAE pre-training strategy.
+S8 Comparison of training iterations
+Table S11: Quantitative details on comparison of finetuning iterations. Our scheme requires much less iterations for finetuning compared to MV-JAR [60].
 Method Data amount Epochs Number of input scans (×$\times 10^{3}$ per epoch) Total iterations (×$\times 10^{5}$) Number of GT used (×$\times 10^{3}$ per epoch) L2 mAPH (Overall) L2 mAPH (Pedestrian)
@@
 20% 48 31.6 15.19 31.6 59.15 59.02
-Random/T-MAE 5% 30 7.9 2.37 4.0 40.29/48.07 44.76/55.71
-10% 30 15.8 4.75 7.9 53.13/57.05 53.04/59.24
-20% 30 31.6 9.50 15.8 57.61/61.72 58.41/64.53
-Table 8: Quantitative details on comparison of finetuning iterations. Our scheme requires much less iterations for finetuning compared to MV-JAR [48].
-Appendix A Additional Results on Waymo Subsets
-The data-efficient benchmark, proposed by MV-JAR [48], contains three splits, each of which contains four tiers of data, namely 5%, 10%, 20% and 50%.
-In the context of data organization, each split corresponds to a subdivision where a lower level of data is a subset of a higher level.
-There is no overlap observed between different splits for the data corresponding to the 5% and 10% subsets.
-In the main paper, the majority of experiments were performed by finetuning the model with the split 0.
-To alleviate the variance caused by being finetuned with a small amount of data, we further present the performance obtained when the model is finetuned using the remaining two splits, as shown in Tab.
-5 and Tab.
-6.
-Furthermore, the average results are presented in Tab.
-7.
-In conclusion, it can be observed that the model performance exhibits variation when finetuned with different subsets.
-However, T-MAE consistently demonstrates a noteworthy improvement in model performance compared to random initialization.
-Additionally, T-MAE attains SOTA average performance across all metrics when finetuned using 10% data.
-Appendix B Visualization of cross-attention scores
-To understand the knowledge acquired by the SiamWCA during pre-training, we employ attention visualization to identify critical regions of the previous scan.
-In particular, the entire previous scan and the target objects of current scan are input into the SiamWCA.
-More precisely, the input to the model for the current scan solely consists of points contained within the 3D ground-truth bounding boxes.
-The purpose of removing other points is to identify the specific areas of emphasis within the WCA module when the queries are solely target objects, e.g.
-vehicles and pedestrians.
-As illustrated in Fig.
-6 (b), the attention is mainly attached to target objects, indicating that the WCA successfully detects and localizes these entities in the previous frame.
-Furthermore, the presence of vehicular action is perceptible even if the WCA module is only trained with unlabeled data, indicating the effectiveness of our T-MAE pre-training strategy.
-Appendix C Comparison of training iterations.
+Random/T-MAE 5% 30 7.9 2.37 4.0 40.29/49.46 44.76/55.28
+10% 30 15.8 4.75 7.9 53.13/57.99 53.04/61.10
+20% 30 31.6 9.50 15.8 57.61/61.80 58.41/64.66
 MV-JAR [60] applies varying numbers of epochs during finetuning, depending on the data amount.
-Rather than varying numbers, we employ a predetermined number of epochs, leading to significantly less iterations for finetuning without performance drop.
+Rather than varying numbers, we employ a predetermined number of epochs, leading to significantly fewer iterations for finetuning without a performance drop.
 Detailed information is presented in Tab.
-8.
+S11.
 The total number of iterations is calculated by multiplying the number of input scans by the number of epochs.
@@
 1 and Tab.
-8, T-MAE achieve better performance than MV-JAR [48] while requiring a smaller number of finetuning iterations.
-In addition, our scheme requires just one GT annotation every two input frames.
-As a result, the total number of iterations would be even smaller if it is calculated by multiplying the number of GT by the number of epochs.
+S11, T-MAE outperforms MV-JAR [60] while requiring a smaller number of finetuning iterations.
+S9 Qualitative Results
+Fig.
+S3 shows the qualitative results of our method on the Waymo dataset [49].
+Fig.
+S4 and Fig.
+S5 illustrates qualitative results on the ONCE datset [37].
+Our method generates accurate bounding boxes even if the scene is complex.
+S10 Limitations
+While our work has achieved encouraging results, there is space for further improvements.
+For instance, the alignment of two frames relies on their transformation matrices; without proper alignment, there is a drop in performance.
+If this could be eliminated, the proposal would be more efficient.
+Another shortcoming is that the transformer blocks are computationally heavy, which restricts the window size.
+This leads to a smaller receptive field and thus makes the network easy to lose track of fast-moving objects.
+The transformer-based encoder also increases the inference time, which was tested on a single NVIDIA RTX 3090 and measured as 429 ms per frame.
+Consequently, the network is not real-time.
+Figure S3: Qualitative results on the Waymo dataset [49]. We depict ground truth and predictions as boxes colored in red and green for several exemplary scenes.
+Figure S4: Qualitative results on the ONCE dataset [37]. We depict ground truth and predictions as boxes colored in red and green for several exemplary scenes.
+Figure S5: Qualitative results on the ONCE dataset [37]. We depict ground truth and predictions as boxes colored in red and green for several exemplary scenes.